### v2

# On the Convergence of Muon and Beyond

Abstract
The Muon optimizer has demonstrated remarkable empirical success in handling matrix-structured parameters for training neural networks. However, a significant gap persists between its practical performance and theoretical understanding. Existing analyses indicate that the standard Muon variant achieves only a suboptimal convergence rate of $\mathcal{O}(T^{-1/4})$ in stochastic non-convex settings, where $T$ denotes the number of iterations. To explore the theoretical limits of the Muon framework, we develop and analyze two momentum-based variance-reduced variants: a one-batch version (Muon-MVR1) and a two-batch version (Muon-MVR2). We provide the first rigorous proof that incorporating a variance-reduction mechanism enables Muon-MVR2 to attain an optimal convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$, thereby matching the theoretical lower bound for this class of problems. Moreover, our analysis establishes convergence guarantees for Muon variants under the Polyak-Łojasiewicz (PŁ) condition. Extensive experiments on vision (CIFAR-10) and language (C4) benchmarks corroborate our theoretical findings on per-iteration convergence. Overall, this work provides the first proof of optimality for a Muon-style optimizer and clarifies the path toward developing more practically efficient, accelerated variants.
1 Introduction
The immense computational cost of pre-training Large Language Models (LLMs) has spurred a surge of research into novel optimization methods designed to enhance parameter efficiency and training stability Hoffmann et al. (2022); Liu et al. (2023); Chen et al. (2023); Vyas et al. (2025); Pethick et al. (2025); Yuan et al. (2024). Among these, methods based on matrix orthogonalization have recently garnered significant attention from both academia and industry Jordan et al. (2024); Liu et al. (2025a). In particular, the Muon optimizer has emerged as a notable milestone due to its impressive empirical performance Liu et al. (2025a); An et al. (2025); Liu et al. (2025b); Shah et al. (2025). However, despite its practical success, the theoretical understanding of Muon’s underlying mechanisms has surprisingly lagged behind, with existing convergence analyses being fraught with limitations and even critical fallacies.
Specifically, the current theoretical exploration of Muon’s convergence faces three primary obstacles. First, some analyses are predicated on flawed mathematical assumptions, such as the erroneous use of the inequality $\|\mathbf{S}^{-1}\|_{2}\leq 1/\|\mathbf{S}\|_{2}$ in a key part of their proof, which casts serious doubt on the validity of their convergence claims Li & Hong (2025). Second, existing analyses have not established convergence to a stationary point without reliance on problem dimension or batch size; the available bounds become valid only when the batch size is sufficiently large, so the basic requirement for a trustworthy optimizer remains unfulfilled Sato et al. (2025); Sfyraki & Wang (2025). Finally, the most rigorous existing work analyzes both the standard Muon algorithm (like Algorithm 1, Option MVR1 with $\gamma=0$) and its simplified variant , considering their convergence to a non-standard $\epsilon$-nuclear norm stationary point. This leaves the behavior of the Nesterov-Accelerated Muon algorithm (like Algorithm 1, Option MVR1), as discussed in Liu et al. (2025a); Sato et al. (2025), and Variance-Reduction Muon (like Algorithm 1, Option MVR2) under standard settings as an open question Shen et al. (2025).
To bridge this critical theoretical gap, this work establishes a rigorous and complete theoretical foundation for the Muon optimizer. Our main contributions are threefold:
  • For general non-convex settings, we provide a convergence analysis for both the standard Muon (Algorithm 1, Option MVR1 with $\gamma=0$) and the Muon-MVR1 algorithm (Algorithm 1, Option MVR1).
  • Regarding the optimal convergence rate, we are the first to demonstrate that the Variance-Reduction Muon-MVR2 algorithm (Algorithm 1, Option MVR2) can overcome the bottleneck of traditional momentum methods, achieving the optimal ergodic convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$.
  • Under the Polyak-Łojasiewicz (PL) condition, we prove that our proposed algorithms, Muon-MVR1, and Muon-MVR2, all achieve sublinear convergence rates. More precisely, we demonstrate that Muon-MVR1 converge at a rate of $\tilde{\mathcal{O}}(1/T^{1/2})$, and Muon-MVR2 converges at an accelerated rate of $\tilde{\mathcal{O}}(1/T^{2/3})$
We also refer readers to see the main contributions of our works and comparisons with the existing works in Table 1.
Table 1: Comparison for existing Muon-type analyses with ours
Smooth
Stoc. Gradient
Estimator
Batch
Size
Iteration
Complexity
Convergence
Rate
Li & Hong (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
Sato et al. (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(n)$
Sato et al. (2025) $L$ Option MVR1 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(n)$
Sfyraki & Wang (2025) $L_{+}$ Option MVR2 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(T^{-1/3})$
Shen et al. (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
Ours $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ Th4.3
Ours $L$ Option MVR1 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ Th4.3
Ours $L_{+}$ Option MVR2 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/3})$ Th4.4
  • a
    Option MVR1 ($\gamma=0$) is the standard momentum method, Option MVR1 is the one-batch variance-reduction momentum method, and Option MVR2 is the two-batch variance-reduction momentum method. We summarize them in Algorithm 1.
  • b
    Although increasing the batch size can mitigate the impact of stochastic noise Sato et al. (2025), these methods still fail to converge to a stationary point and cannot eliminate the influence of dimensionality.
  • c
    The results from Sfyraki & Wang (2025) on Option MVR2 are the closest to ours. However, their method requires a large initial batch of size $\mathcal{O}(\epsilon^{-1})$, although the batch size can be reduced to 1 in subsequent iterations.
    $L$ pertains to the smoothness of the overall function $f$, while $L_{+}$ pertains to the smoothness of its stochastic components $f(\cdot;\xi)$.
2 Preliminaries and Problem Setup
2.1 Notations
We denote scalars by non-bold letters (e.g., $a,A$), vectors in $\mathbb{R}^{d}$ by bold lowercase letters (e.g., $\mathbf{a}$), and matrices by bold uppercase letters (e.g., $\mathbf{A}$). The space $\mathbb{R}^{d}$ is endowed with the Euclidean inner product $\langle\mathbf{x},\mathbf{y}\rangle_{2}:=\mathbf{x}^{\top}\mathbf{y}$ and norm $\|\mathbf{x}\|_{2}$. For matrices, we employ the Frobenius inner product $\langle\mathbf{A},\mathbf{B}\rangle_{F}:=\text{tr}(\mathbf{A}^{\top}\mathbf{B})$ and the corresponding norm $\|\mathbf{A}\|_{F}$. The nuclear norm, denoted by $\|\mathbf{A}\|_{*}$, is defined as the sum of the singular values of the matrix, $\|\mathbf{A}\|_{*}=\sum_{i}\sigma_{i}(\mathbf{A})$. Throughout the paper, $[m]$ denotes the set of integers $\{1,2,\ldots,m\}$, and $\mathbb{N}$ denotes the set of non-negative integers.
The model is parameterized by a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$. Without loss of generality, we assume $m\geq n$, so the rank of the matrix is at most $n$. The model is optimized by minimizing the empirical loss function $f(\mathbf{X}):=\frac{1}{N}\sum_{i\in[N]}f_{i}(\mathbf{X})$, where $N$ is the number of training data points, and $f_{i}(\mathbf{X})$ is the loss function for $\mathbf{X}\in\mathbb{R}^{m\times n}$ with respect to the $i$-th training data point $\mathbf{z}_{i}$ (for $i\in[N]$). Let $\xi$ be a random variable that is independent of $\mathbf{X}\in\mathbb{R}^{m\times n}$, and let $\mathbb{E}_{\xi}[\mathbf{X}]$ denote the expectation of a random variable $\mathbf{X}$ with respect to $\xi$.
2.2 Problem Setup
We consider minimizing the objective function as follows
$\min_{\mathbf{X}}f(\mathbf{X}),\mathrm{~where~}f(\mathbf{X})=\mathbb{E}_{\xi\sim\mathcal{D}}[f(\mathbf{X};\xi)].$ (1)
For problem (1), let $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ be a function where $\mathbf{X}\in\mathbb{R}^{m\times n}$ and $\xi$ represents a random variable, such as a training data point, sampled from an unknown data distribution $\mathcal{D}$. We assume that $f$ is differentiable and possibly nonconvex.
3 Revisiting the Muon Algorithms
Algorithm 1 Muon
1:Require Initial parameters $\mathbf{X}_{0}\in\mathbb{R}^{m\times n}$, learning rate $\eta_{t}>0$, momentum parameter $\beta_{t}\in[0,1)$, variance-reduction parameter $\gamma\in[0,1]$, initial momentum $\mathbf{M}_{0}=0$, $\nabla f(\mathbf{X}_{0};\xi)=0$.
2:for $t=1$ to $T$ do
3:  Compute stochastic gradient: $\nabla f(\mathbf{X}_{t};\xi_{t})$
4:  Option MVR1: One-batch Momentum Variance-Reduction (MVR1)
5:  $\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1}))$
6:  Option MVR2: Two-batch Momentum Variance-Reduction (MVR2)
7:  $\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t}))$
8:  $\mathbf{O}_{t}\in\arg\min_{\mathbf{O}}\|\mathbf{O}-\mathbf{M}_{t}\|_{F},\text{ s.t. }\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}_{n}$
9:  $\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}\mathbf{O}_{t}$
10:end for
11:Return Final parameters $\mathbf{X}_{t}$
In practice, Algorithm 1 leverages a quintic Newton–Schulz iteration to efficiently compute a rank-preserving normalization of any input matrix $\mathbf{M}_{t}\in\mathbb{R}^{m\times n}$. This operation can be precisely defined via the Compact Singular Value Decomposition (SVD) of $\mathbf{M}_{t}$, expressed as $\mathbf{M}_{t}=\mathbf{U}_{r}\mathbf{\Sigma}_{r}\mathbf{V}_{r}^{\top}$, where $r$ is the rank of $\mathbf{M}_{t}$ and $\mathbf{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ contains only the $r$ positive singular values. The operator computes the matrix $\mathbf{O}_{t}\in\mathbb{R}^{m\times n}$ given by:
$\mathbf{O}_{t}=\mathbf{U}_{r}\mathbf{V}_{r}^{\top}$
This is equivalent to setting all positive singular values of $\mathbf{M}_{t}$ to 1 while leaving its null space unchanged, thus preserving the original rank $r$. It is worth noting that this formulation is algebraically equivalent to using the Moore-Penrose pseudoinverse: $((\mathbf{M}_{t}\mathbf{M}_{t}^{\top})^{+})^{1/2}\mathbf{M}_{t}$.
Only in the specific case where $\mathbf{M}_{t}$ is full-rank does this result coincide with the solution to the Orthogonal Procrustes problem, $\min_{\mathbf{O}}\|\mathbf{O}-\mathbf{M}_{t}\|_{F}^{2}$ s.t. $\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}$. In the subsequent theoretical analysis, we assume that this operator can be solved exactly. Building upon this core operator, we investigate the following update rules:
[…]
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1}))$ (2)
Further analysis reveals that when the parameters are set as $\beta_{t}=\mu$ and $\gamma=1-\beta_{t}$, the iterative process of Algorithm 1, subject to a rescaling of $\mathbf{M}_{t}$, becomes mathematically equivalent to the update scheme (3) of the Muon algorithm Jordan et al. (2024); Liu et al. (2025a). The Muon update is given by: $\mathbf{M}_{t}=\mu\mathbf{M}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t})+\mu(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1}))$
$\mathbf{C}_{t}$ $=\mu\mathbf{C}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t})$ (3)
$\mathbf{M}_{t}$ $=\mu\mathbf{C}_{t}+\nabla f(\mathbf{X}_{t};\xi_{t})$
$\mathbf{O}_{t}$ $=\text{Newton-Schulz}(\mathbf{M}_{t})$
$\mathbf{X}_{t+1}$ $=\mathbf{X}_{t}-\eta\mathbf{O}_{t}$
Muon-MVR1 reduces to the EMA stochastic gradient estimator Kingma & Ba (2014) update Rule (4) as a special case when $\gamma=0$, and it can also be regarded as an approximation of MVR2.
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})$ (4)
[…]
$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t}))$ (5)
The key distinction between this rule and Rule (2) lies in the term subtracted from the current stochastic gradient, $\nabla f(\mathbf{X}_{t};\xi_{t})$. Specifically, Rule (5) subtracts the previous gradient computed using the current data sample $\xi_{t}$, namely $\nabla f(\mathbf{X}_{t-1};\xi_{t})$, a crucial modification designed to cancel the variance introduced by $\xi_{t}$ Yuan et al. (2024). In contrast, Rule (2) reuses the gradient based on the previous data sample $\xi_{t-1}$ (i.e., $\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$), effectively acting as a computationally efficient approximation of Rule (5).
This approximation is justified because the difference between the two approaches is a noise term, $\Delta^{\text{Noise}}_{t-1}=\nabla f(\mathbf{X}_{t-1};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$. Under the standard assumption that stochastic gradients have a bounded variance, $\mathbb{E}_{\xi}[\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}]\leq\sigma^{2}$, the variance of this noise term is well-controlled and satisfies $\mathbb{E}[\|\Delta^{\text{Noise}}_{t-1}\|_{F}^{2}]\leq 2\sigma^{2}$. While Rule (2) offers the practical benefit of requiring only one gradient evaluation per step versus two for Rule (5), this structural difference leads to fundamentally different convergence properties. Our theoretical analysis demonstrates that the Muon-variant algorithm employing the stricter variance-reduction mechanism of Rule (5) achieves a convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ in Theorem 4.2.
4 Convergence Analysis
We begin our analysis by situating it within the established context of first-order stochastic optimization. The $\mathcal{O}(T^{-1/4})$ convergence rate is a well-known bottleneck for methods like SGD, and recent work by Shen et al. (2025) confirmed this limitation for the standard Muon algorithm. Accordingly, our first step in this section is to analyze the Muon-MVR1 variant and formally reaffirm this baseline convergence rate in Theorem 4.1. While the rate itself is standard, our primary contribution here lies in the novel and unified analytical framework we develop to prove it, as this framework will be instrumental for subsequent results.
Our analysis is conducted under a diminishing stepsize schedule, a standard setting that guarantees convergence to an optimal solution. Nevertheless, our theoretical framework is general enough to encompass the constant-stepsize setting, where it yields a tighter convergence bound devoid of the logarithmic factor $\ln(T)$.
To facilitate the analysis of convergence for Muon, we make the following assumptions:
Assumption 4.1.
The function $f$ is bounded from below. There exists $f^{*}>$-$\infty$ such that $f(\mathbf{X})\geq f^{*}$, for all $\mathbf{X}\in\mathbb{R}^{m\times n}$.
Assumption 4.2.
The function $f$ is $L$-smooth: $\|\nabla f(\mathbf{Y})-\nabla f(\mathbf{X})\|_{F}\leq L\|\mathbf{Y}-\mathbf{X}\|_{F}$.
Assumption 4.3.
The function $f$ is $L$-smooth for any $\xi$: $\|\nabla f(\mathbf{Y};\xi)-\nabla f(\mathbf{X};\xi)\|_{F}\leq L\|\mathbf{Y}-\mathbf{X}\|_{F}$.
Assumption 4.4.
The variance of unbiased stochastic gradient is finite. Specifically, there exists a constant $\sigma>0$ such that for all $\mathbf{X}\in\mathbb{R}^{m\times n}$, the following holds: $\mathbb{E}[\nabla f(\mathbf{X};\xi)]=\nabla f(\mathbf{X})$ and $\mathbb{E}\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}\leq\sigma^{2}$.
These assumptions are quite common Zhou et al. (2018); Chen et al. (2018); Huang et al. (2021); Guo et al. (2021); Li et al. (2023); Wang et al. (2023); Xie et al. (2024); Yuan et al. (2024).
4.1 Erogdic Convergence of Muon
4.1.1 Option MVR1
Theorem 4.1.
Suppose Assumptions 4.1, 4.2, and 4.4 hold. Consider Algorithm 1 with a learning rate of $\eta_{t}=t^{-3/4}$. The expected average squared norm of the gradient is bounded for the following options:
  1. 1.
    For Option MVR1 ($\gamma=0$), with momentum parameter $\beta_{t}=1-t^{-1/2}$, the bound is given by:
    $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{*}}{T^{1/4}}+\frac{A_{1}\ln T+A_{2}}{T^{1/4}},$
    where $A_{1}=2L^{-1}\sigma^{2}+4\sqrt{2}Ln+Ln+L/2$ and $A_{2}=4L^{-1}\sigma^{2}+4\sqrt{2}Ln+Ln+L/2$.
  2. 2.
    For Option MVR1 ($\gamma_{t}=t^{-1/2}$), with momentum parameter $\beta_{t}=1-(t+1)^{-1/2}$, the bound is given by:
    $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{*}}{T^{1/4}}+\frac{A_{1}\ln T+A_{2}}{T^{1/4}},$
    where $A_{1}=4L^{-1}\sigma^{2}+8\sqrt{2}Ln+Ln+L/2$ and $A_{2}=10L^{-1}\sigma^{2}+8\sqrt{2}Ln+Ln+L/2$.
See Appendix B for details.
Remark 4.1.
Theorems 4.1 establishes that as $T\to\infty$, the leading terms diminish to zero, and the algorithm converges to a neighborhood of a stationary point. The size of this neighborhood is determined by the stochastic gradient variance, the learning rate, and the momentum parameter. Notably, this convergence bound is free of any non-vanishing additive error term that depends on the dimension $n$ Sato et al. (2025). We absorb the dimensional dependence into the $\mathcal{O}$-notation to define $\tilde{\mathcal{O}}$, which yields the following complexity for the algorithm:
$\min_{t=1,\dots,T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\tilde{\mathcal{O}}(T^{-1/4})$
4.1.2 Option MVR2
While prior work has established a convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ for variance-reduction in non-convex settings for various algorithms, such as SGD Fang et al. (2018); Cutkosky & Orabona (2019); Zhou et al. (2020) and Adam Huang et al. (2021); Yuan et al. (2024), the theoretical convergence properties of this technique when applied to the Muon optimizer have remained an open question. We bridge this theoretical gap in Theorem 4.2, which rigorously proves that Option MVR2 of Algorithm 1 (Muon-MVR2) achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ convergence rate in the general non-convex setting.
Theorem 4.2.
Under Assumptions 4.1, 4.3, and 4.4, consider Algorithm 1 with Option MVR2 and update rule (5) parameters set as $\beta_{t}=1-\eta_{t}$, $\eta_{t}=t^{-2/3}$ for $t\geq 1$, and $\gamma=1$. Then, the following bounds hold:
[…]
where
$G=f(\mathbf{X}_{1})-f^{*}+2\sigma^{2}L^{-1}+2Ln+\left(8Ln+4\sigma^{2}L^{-1}+\frac{Ln}{2}\right)(1+\ln T).$
See Appendix D for details.
Remark 4.2.
Theorem 4.2 establishes a significant theoretical advantage. In contrast to prior work Sfyraki & Wang (2025) that requires a computationally expensive growing batch size ($b=\mathcal{O}(T^{1/3})$) to control gradient variance, our result shows that Muon-MVR2 achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ convergence rate in a more practical, fixed mini-batch setting. This is made possible by the specific hyperparameter schedule, where the learning rate $\eta_{t}=t^{-2/3}$ and the momentum parameter $\beta_{t}=1-\eta_{t}$ are critically coupled. This schedule ensures a careful balance between optimization progress and noise suppression.
Remark 4.3.
As noted in Yuan et al. (2024), a more sophisticated, adaptive setting for the variance-reduction parameter $\gamma$ can be employed. Specifically, by setting $\gamma=\gamma_{t}=1-\frac{A_{t}}{\beta_{t}}$, where $A_{t}$ is defined in Lemma C.2, a key term in the analysis becomes $P_{t}=-\mathbb{E}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t})\|_{F}^{2}\cdot A_{t}^{2}$, which is strictly negative. This leads to a tighter convergence bound, as it effectively introduces an additional beneficial term into the recurrence. However, computing this adaptive $\gamma_{t}$ is often impractical as it depends on quantities that are difficult to estimate during training. Consequently, we adhere to the common and more practical approach of using a constant $\gamma\leq 1$.
4.2 Non-Ergodic Convergence of Muon
In this subsection, we examine the performance of Muon-MVR1 and Muon-MVR2 under the setting where the non-convex objective functions satisfy the Polyak-Łojasiewicz (PL) condition. Our analysis is based on the following additional assumption:
Assumption 4.5.
We assume the function $f$ is $\mu$-PL, i.e., $\|\nabla f(\mathbf{X}_{t})\|_{F}^{2}\geq 2\mu(f(\mathbf{X}_{t})-f^{*})$.
Remark 4.4.
The PL condition has been widely employed in the convergence analysis of various first-order algorithms Karimi et al. (2016); Xie et al. (2020a); Li & Li (2022), though typically under restricted settings. Note that if $f(\cdot)$ is strongly convex, then it is necessarily convex and satisfies the PL condition. However, the converse does not hold in general; a counterexample is given by $f(x)=x^{2}+3\sin(2x)$.
Theorem 4.3.
Suppose Assumptions 4.1, 4.2, 4.4, and 4.5 hold. Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with a step size of $\eta_{t}=t^{-3/4}$. For either of the following options:
  1. 1.
    For Option MVR1 ($\gamma=0$), using the update rule (4) with momentum $\beta_{t}=1-t^{-1/2}$,
  2. 2.
    For Option MVR1 ($\gamma_{t}=t^{-1/2}$), using the update rule (2) with parameters $\beta_{t}=1-(t+1)^{-1/2}$,
the expected suboptimality gap after $T$ iterations is bounded by:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right).$
Theorem 4.4.
Suppose Assumptions 4.1, 4.2, 4.4, and 4.5 hold. Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with Option MVR2 (see update rule (5)). By setting the parameters to $\eta_{t}=t^{-2/3}$, $\beta_{t}=1-t^{-2/3}$, and $\gamma=1$, the expected suboptimality gap after $T$ iterations is bounded by:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right).$
Remark 4.5.
The convergence bounds in Theorems 4.3, and 4.4 are asymptotic, holding for any number of iterations $T$ exceeding a sufficiently large threshold $T_{A}$. The existence of this threshold is established in our proofs and depends on the selection of a helper constant $A$. The specific conditions on $A$ and definitions of $\mathcal{A}$ are as follows:
  • For Theorem 4.3 (MVR1 with $\gamma=0$): $A>\frac{\mathcal{A}}{8C}$ with $\mathcal{A}=6L^{-1}\sigma^{2}+8\sqrt{2}Ln+2Ln+L$.
  • For Theorem 4.3 (MVR1): $A>\frac{\mathcal{A}}{8C}$ with $\mathcal{A}=14L^{-1}\sigma^{2}+16\sqrt{2}Ln+2Ln+L$.
  • For Theorem 4.4 (MVR2): $A>\frac{\mathcal{A}}{6C}$ with $\mathcal{A}=20L^{-1}\sigma^{2}+34Ln+L$.
In all cases, the constant $C$ is defined as $C=\frac{\sqrt{2\mu}}{2}$. The proofs of Theorems 4.3, and 4.4 are provided in Appendix F, and G, respectively.
Remark 4.6.
Theorem 4.4 achieves a faster convergence rate of $\tilde{\mathcal{O}}(T^{-2/3})$ than the $\tilde{\mathcal{O}}(T^{-1/2})$ in Theorem 4.3, owing to its more effective variance reduction. While our rates are not as fast as the $\mathcal{O}(1/T)$ rate for Adam He et al. (2023), our analysis is more general, as it does not require the bounded stochastic gradient assumption used therein. Closing this rate gap under our relaxed assumption is an important direction for future work.
Remark 4.7.
Our non-ergodic convergence proofs are unified by a key recursive inequality for a Lyapunov function $\Delta_{t}=\mathbb{E}[f(\mathbf{X}_{t})]-f^{*}$:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}.$
The final convergence rate is dictated by the exponent $p\in(0,1)$. We establish this with $p=3/4$ for Theorem 4.3. For Theorem 4.4, superior variance control allows proving the inequality with a tighter $p=2/3$, leading to the faster rate.
5 Experiments
In this section, we evaluate the performance of the Muon-variant optimizers on pretraining tasks. All experiments were conducted using four Ascend 910C (64GB) NPUs and four NVIDIA RTX 4090 (24GB) GPUs. Detailed experimental settings are provided in Appendix H.
(a) Train and test accuracy curves
(b) Train and test loss curves
(c) Test accuracy over wall-clock time
Figure 1: Training dynamics of Muon-MVR2, Muon-MVR1, Muon-MVR1 ($\gamma=0$), and a baseline on CIFAR-10 with ResNet-18. The plots show (a) accuracy and (b) loss versus epochs for both training and testing, along with (c) test accuracy versus wall-clock time.
[…]
(a) Train loss curve
(b) Validation loss curve
Figure 2: LLaMA2-130M train and validation curves on C4 Dataset
Remark 5.1.
The precise, Variance-Reduction formulation (5) is more rigorously aligned with stochastic optimization theory and, as demonstrated in our experiments, consistently achieves the highest validation accuracy. However, its computational overhead can be prohibitive in large-scale settings, as shown in Figure 1c. We note that prior analysis in Sfyraki & Wang (2025); Sato et al. (2025) suggests that increasing batch size can mitigate the impact of stochastic noise on convergence properties. Given that mini-batch training is the standard paradigm in practice, a trade-off between theoretical exactness and computational feasibility is often inherent. Therefore, echoing the sentiment in Yuan et al. (2024), we posit that the approximate version serves as a pragmatic alternative when computational efficiency is a primary concern, as it incurs only a minimal degradation in performance. Conversely, in scenarios where maximizing validation accuracy is paramount, the precise formulation remains the recommended approach.
6 Related Work
6.1 The Evolution of Optimization Algorithms
One of the cornerstones of deep learning is first-order optimization methods. In early research, the Momentum method proposed by Nesterov (1983) accelerates the convergence process by accumulating historical gradient information. Subsequently, the RMSprop algorithm introduced by Hinton et al. (2012) pioneered the concept of adaptive learning rates, allowing different parameters to have independent update step sizes. The Adam optimizer, proposed by Kingma & Ba (2014), integrated these two ideas and has become one of the most widely used optimizers in the deep learning field by adaptively adjusting the first and second moments of the gradients.
Building upon Adam, the academic community has proposed a series of improvements to address its limitations in various scenarios. To enhance the stability of the optimization process, Reddi et al. (2018) proposed AMSGrad, which uses the maximum of historical second moments to constrain the learning rate update. Dozat (2016) integrated Nesterov momentum into Adam to design NAdam, aiming for superior performance. To address issues in Adam concerning weight decay and learning rate adjustment, Loshchilov (2017) developed AdamW, which improves the effectiveness of L2 regularization by decoupling weight decay. Luo et al. (2019) designed AdaBound, which sets dynamic bounds on the learning rate to avoid excessively large or small update steps. Furthermore, to mitigate variance fluctuations in the early stages of training, Liu et al. (2019) proposed RAdam to rectify the variance estimate. Zhuang et al. (2020) proposed AdaBelief, which adjusts the second moment by calculating the deviation between the gradient and its moving average, aiming to enhance the model’s generalization ability. In recent years, more novel Adam variants have been proposed. For instance, Xie et al. (2020b), from a dynamical systems perspective, constructed the Adai framework to accelerate training and improve the quality of convergence points by decoupling the effects of the adaptive learning rate and momentum. Xie et al. (2024) proposed Adan, which introduces a novel Nesterov momentum estimation method to accelerate convergence without incurring additional gradient computation overhead. The latest works, such as Liang et al. (2024), introduce a Cautious masking strategy to enhance optimization efficiency. Yuan et al. (2024) proposed MARS to further improve the variance-reduction strategy in Adam.
Beyond direct improvements to Adam, researchers have also explored other efficient optimization paradigms. A significant branch among these is Shampoo-type methods based on preconditioning with parameter curvature. Shampoo, proposed by Gupta et al. (2018), is one of the pioneering works in this area. Inspired by its framework, Muon Jordan et al. (2024); Liu et al. (2025a) was proposed. This method adapts to parameter curvature by orthogonalizing gradient momentum via the Newton-Schulz iteration, thereby improving convergence dynamics. Recently, several adaptive variants of Muon have also emerged. AdaMuon Si et al. (2025) incorporates element-wise adaptivity but lacks theoretical convergence guarantees. For efficient training of large language models, COSMOS Liu et al. (2025b) integrates ideas from SOAP Vyas et al. (2025) and Muon, demonstrating stability and memory advantages in practice but similarly lacks a convergence proof. In the literature, other related preconditioning methods include ASGO An et al. (2025), PolarGrad Lau et al. (2025), and AdaGO Zhang et al. (2025), which introduces the Adagrad-Norm step size Ward et al. (2020) into a simplified version of Muon. Meanwhile, other high-performing optimizers not belonging to the Shampoo family also warrant attention, such as Sophia Liu et al. (2023), which improves second-moment estimation through efficient diagonal Hessian approximation and coordinate clipping, and Lion Chen et al. (2023), a lightweight optimizer that only tracks momentum and uses the sign function to normalize updates.
6.2 Analysis of Muon
Recent studies have elucidated the underlying mechanisms and convergence properties of the Muon optimizer from multiple theoretical perspectives. A central viewpoint reinterprets modern optimizers as steepest descent methods under norm constraints. Although this analysis does not directly incorporate momentum, it offers valuable insights into the update directions of optimizers like Muon and Adam Bernstein & Newhouse (2024). Building on this, subsequent research has established a theoretical link between the Muon algorithm with weight decay and the stochastic Frank-Wolfe method, specifically identifying Muon as an instance of stochastic Frank-Wolfe under a spectral norm constraint Sfyraki & Wang (2025). This provides a foundational understanding of its behavior from a constrained optimization perspective. Other work has characterized Muon as a special case of Lion-$\mathcal{K}$. Separately, its convergence to a Karush-Kuhn-Tucker (KKT) point has been proven in both deterministic and stochastic gradient settings Chen et al. (2025). These theories, together with research on stochastic conditional gradient methods from a norm-constrained viewpoint, form the theoretical underpinnings for Muon Pethick et al. (2025). Beyond its theoretical framework, Muon’s implicit bias has garnered significant attention. It has been shown that Muon tends to converge to a max-margin solution with respect to the spectral norm of the classifier’s weight matrix, highlighting a distinct difference in implicit regularization and generalization preference compared to Adam Fan et al. (2025). A unifying preconditioning perspective explains this phenomenon through matrix polar decomposition, which explicitly distinguishes between curvature anisotropy and gradient anisotropy in optimization problems, thereby clarifying the behavioral differences between Muon and Adam when handling different types of parameters Lau et al. (2025). Regarding convergence analysis, while rigorous proofs and convergence rates for Muon in non-convex settings have been provided under various smoothness assumptions (e.g., Frobenius-norm or spectral-norm Lipschitz smoothness) Shen et al. (2025), the validity of some early analyses has been questioned. For instance, certain convergence results are considered dubious due to the use of incorrect mathematical inequalities in their proofs Li & Hong (2025). Furthermore, other analyses, despite proving the convergence of several Muon variants, have indicated slow convergence rates or the need for stringent conditions to reach a desired stationary point Sato et al. (2025). On a practical level, particularly in the pre-training of large-scale language models (LLMs), weight decay has been shown to be an indispensable component. Concurrently, methods for effectively transferring learning rates from Adam to Muon, based on RMS analysis, have been proposed to enhance its scalability and utility Liu et al. (2025a).
7 Conclusion
In this work, we established a rigorous theoretical foundation for the Muon optimizer, addressing the significant gap between its empirical success and its formal analysis. We introduced Muon-MVR2, a variance-reduced variant, and provided the first proof that it achieves an optimal convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ in stochastic non-convex settings, matching the theoretical lower bound Arjevani et al. (2023). Furthermore, we demonstrated that under the Polyak-Łojasiewicz condition, the Muon-MVR1, and Muon-MVR2 variants all attain sublinear non-ergodic convergence. Our theoretical results are validated by extensive experiments on CIFAR-10 and C4 benchmarks, which confirm the practical acceleration and superior performance of Muon-MVR2 over its standard counterpart and other widely-used optimizers. This research solidifies the Muon framework by providing robust theoretical guarantees and a practically effective new variant for training deep learning models.
Appendix
Appendix A Lemmas for Theorem 4.1
A.1 Lemma A.1
Lemma A.1.
For Algorithm 1, choosing an arbitrary parameter $\alpha>0$, we have the following inequality:
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta^{2}n}{2}$
Proof.
According to Assumption 4.2, we have the upper bound for the function value:
$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})+\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{t})-\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{t})-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle-\eta_{t}\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
Now we bound the three terms on the right-hand side respectively:
Descent Term: According to the definition of $\mathbf{O}_{t}$ and the norm property $\|\mathbf{M}_{t}\|_{*}\geq\|\mathbf{M}_{t}\|_{F}$, we have:
$\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}$
Cross Term: This is the key to eliminating the dimension-dependent error. We use Young’s inequality with a parameter ($ab\leq\frac{\alpha}{2}a^{2}+\frac{1}{2\alpha}b^{2}$) and the fact that $\|\mathbf{O}_{t}\|_{F}^{2}\leq\sum_{i=1}^{n}1=n$:
$-\eta_{t}\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle$ $\leq\eta_{t}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}\|\mathbf{O}_{t}\|_{F}$
$\leq\eta_{t}\left(\frac{\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{1}{2\alpha}\|\mathbf{O}_{t}\|_{F}^{2}\right)$
$\leq\eta_{t}\left(\frac{\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{n}{2\alpha}\right)$
Quadratic Term:
$\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}\leq\frac{L\eta_{t}^{2}n}{2}$
Substituting these three bounds into the inequality for $f(\mathbf{X}_{t+1})$:
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta_{t}^{2}n}{2}$
This completes the proof. ∎
A.2 Lemma A.2
Lemma A.2.
For Algorithm 1 option MVR1 ($\gamma=0$), the accumulated error between the momentum term and the true gradient is bounded:
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}$
Proof.
First, we have
$\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$=\|\beta_{t+1}\mathbf{M}_{t}+(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$=\|\beta_{t+1}(\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}))+(1-\beta_{t+1})(\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1}))$
$\quad+\beta_{t+1}(\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}))\|_{F}^{2}$
$=\beta_{t+1}^{2}\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+\beta_{t+1}^{2}\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$\quad+(1-\beta_{t+1})^{2}\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$\quad+2\beta_{t+1}^{2}\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
According to Assumption 4.4. Taking the expectation of its squared norm, and using the unbiasedness and independence of the stochastic gradient, we obtain:
$\mathbb{E}[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}]=$ $\beta_{t+1}^{2}\mathbb{E}[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}]$
$\quad+(1-\beta_{t+1})^{2}\mathbb{E}[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}]$
$+\beta_{t+1}^{2}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}]$
$+2\beta_{t+1}^{2}\mathbb{E}[\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle]$
Applying Young’s inequality with a parameter ($ab\leq\frac{\epsilon}{2}a^{2}+\frac{1}{2\epsilon}b^{2}$), we have
$\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}\leq\frac{\epsilon}{2}\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+\frac{1}{2\epsilon}\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}.$
Thus, we have:
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$ $\leq\beta_{t+1}^{2}(1+\epsilon)\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]$
$\quad+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$
$\quad+(1-\beta_{t+1})^{2}\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$
According to Assumption 4.2,
$\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$ $\leq L^{2}\|\mathbf{X}_{t}-\mathbf{X}_{t+1}\|_{F}^{2}$
$=L^{2}\eta_{t}^{2}\|\mathbf{O}_{t}\|_{F}^{2}$
$\leq L^{2}\eta_{t}^{2}n$
Therefore:
$\mathbb{E}\left[\left\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\right\|_{F}^{2}\right]$ $\leq\beta_{t+1}^{2}(1+\epsilon)\mathbb{E}\left[\left\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\right\|_{F}^{2}\right]$
$+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
Then, by letting $\epsilon:=\frac{1-\beta_{t}}{\beta_{t}}$, we have
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}t\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}$ (6)
A.3 Lemma A.3
Lemma A.3.
Suppose that $\{E_{i},A_{i}\}$ are two nonnegative sequences. Assume $E_{t+1}\leq(1-\alpha_{t+1})E_{t}+A_{t+1}$ where $\alpha_{t}=t^{-p}$, $p\in(0,1]$. Then we have:
$\alpha_{t}E_{t}\leq 2(E_{t}-E_{t+1}+A_{t+1}).$
Proof.
$\alpha_{t}E_{t}-c\left(E_{t}-E_{t+1}+A_{t+1}\right)$
$\stackrel{{\scriptstyle(\bullet)}}{{\leq}}\alpha_{t}E_{t}-c\left(E_{t}+A_{t+1}\right)+c\cdot\left(E_{t}-\alpha_{t+1}E_{t}+A_{t+1}\right)$
$=E_{t}\left(\alpha_{t}-c\alpha_{t+1}\right)$
$=E_{t}\cdot(t+1)^{-p}\cdot\left(\left(\frac{t}{t+1}\right)^{-p}-c\right)$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}E_{t}\cdot(t+1)^{-p}\cdot(2-c),$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}0$
where $(\bullet)$ follows from $E_{t+1}\leq(1-\alpha_{t+1})E_{t}+A_{t+1}$; $(\circ)$ is due to $(\frac{t}{t+1})^{-p}\leq 2^{p}\leq 2$; $(\star)$ is due to our choice $c=2$. ∎
Appendix B Proofs of Theorem 4.1
Proof.
According to Lemma A.1, we have:
$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta_{t}^{2}n}{2}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}^{2/3}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}^{4/3}Ln}{2}+\frac{L\eta_{t}^{2}n}{2}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}^{2/3}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+Ln\eta_{t}^{4/3}$
where $(\circ)$ by setting $\alpha=\frac{1}{\eta_{t}^{1/3}L}$; $(\star)$ follows from $\eta_{t}\leq 1$, we have $L\leq L/\eta_{t}^{1/3}$.
Thus, taking the expectation yields
$\mathbb{E}[f(\mathbf{X}_{t+1})]$ $\leq\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\mathbf{M}_{t}\|_{F}]+\frac{\eta_{t}^{2/3}}{2L}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}$ (7)
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]+\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}]$
$+\frac{\eta_{t}^{2/3}}{2L}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]+\frac{1}{2\epsilon}\eta_{t}^{2}+\frac{\epsilon}{2}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]$
$+\frac{\eta_{t}^{2/3}}{2L}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}$
$\stackrel{{\scriptstyle(\bullet)}}{{=}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]$
$+\underbrace{\frac{L\eta_{t}^{4/3}}{2}+\left(\frac{\eta_{t}^{2/3}}{2L}+\frac{\eta_{t}^{2/3}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}}_{\Gamma_{t}}$
where $(\circ)$ follows from the reverse triangle inequality $\|\mathbf{M}_{t}\|_{F}\leq\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}-\|\nabla f(\mathbf{X}_{t})\|_{F}$; $(\star)$ applies Young’s inequality to the term $\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}]$; and $(\bullet)$ collects the residual terms into $\Gamma_{t}$ and sets $\epsilon=\eta_{t}^{2/3}/L$.
Next, we set $\eta_{t}=t^{-3/4}$ and $\beta_{t}=1-t^{-1/2},\alpha_{t}=t^{-1/2}$.
Case 1: $\gamma=0$.
By Lemma A.2 inequality (6), we have
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$ $\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}$
$\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{L^{2}\eta_{t}^{2}n}{1-\beta_{t+1}}+(1-\beta_{t+1})^{2}\sigma^{2}$
Let $\mathbf{S}_{t+1}=\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})$. Thus, setting $\alpha_{t+1}=(t+1)^{-1/2}$, we observe the following relationship for $t\geq 1$:
$\frac{\eta_{t}^{2}}{1-\beta_{t+1}}=\frac{t^{-3/2}}{(t+1)^{-1/2}}=\frac{\sqrt{t+1}}{t^{3/2}}\leq\frac{2\sqrt{2}}{t+1}$
This allows us to bound the expectation as follows:
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\alpha_{t+1}^{2}(2\sqrt{2}L^{2}n+\sigma^{2})$
According to Lemma A.3, by letting $A_{t+1}=\alpha_{t+1}^{2}(2\sqrt{2}L^{2}n+\sigma^{2})$, we have
$\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1})$
Furthermore, since
$\mathbb{E}\|\mathbf{S}_{1}\|$ $=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\mathbf{M}_{1}\|_{F}^{2}=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-(1-\beta_{1})\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}$
$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{1}=0)$
It follows that
$\sum_{t=1}^{T}\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq 2\sum_{t=1}^{T}(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1})$ (8)
$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t+1}$
$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
Thus,
$\Gamma_{t}$ $=\frac{\eta_{t}^{2/3}}{L}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\eta_{t}^{4/3}$
$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}$
Next, we define $A_{1}=4\sqrt{2}Ln+Ln+2L^{-1}\sigma^{2}+L/2$ and $A_{2}=4\sqrt{2}Ln+Ln+4L^{-1}\sigma^{2}+L/2$.
$\sum_{t=1}^{T}\Gamma_{t}$ $=\frac{1}{L}\sum_{t=1}^{T}\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\sum_{t=1}^{T}\alpha_{t}^{2}$ (9)
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)}{L}+(L/2+Ln)(\ln T+1)$
$\leq A_{1}\ln T+A_{2}$
where $(\circ)$ due to inequality (8).
Therefore, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $=\frac{1}{T}\sum_{t=1}^{T}t^{3/4}\cdot t^{-3/4}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$
$\leq\frac{1}{T}\sum_{t=1}^{T}t^{3/4}\cdot\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$
$\leq\frac{T^{3/4}}{T}\sum_{t=1}^{T}f(\mathbf{X}_{t})-f(\mathbf{X}_{t+1})+\frac{T^{3/4}}{T}\sum_{t=1}^{T}\Gamma_{t}$
$\leq\frac{f(\mathbf{X}_{1})-f^{*}}{T^{1/4}}+\frac{A_{1}\ln T+A_{2}}{T^{1/4}}$
$=\tilde{\mathcal{O}}(T^{-1/4})$
Case 2: $\gamma\neq 0$.
We set $\eta_{t}=t^{-3/4}$ and $\beta_{t}=1-(t+1)^{-1/2},\alpha_{t}=t^{-1/2}$, and $\gamma=1-\beta_{t-1}=t^{-1/2}$, we first note that an equivalent form of Algorithm 1 Option MVR1 is given by
$\mathbf{C}_{t}$ $=\beta_{t-1}\mathbf{C}_{t-1}+(1-\beta_{t-1})\nabla f(\mathbf{X}_{t};\xi_{t})$ (10)
$\mathbf{M}_{t}$ $=\beta_{t}\mathbf{C}_{t}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})$
In this case, then we have
$\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}$
$=\mathbb{E}\|\beta_{t-1}\mathbf{C}_{t-1}+(1-\beta_{t-1})\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}^{2}$
$=\mathbb{E}\|\beta_{t-1}(\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1}))+(1-\beta_{t-1})(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t}))$
$\quad+\beta_{t-1}(\nabla f(\mathbf{X}_{t-1})-\nabla f(\mathbf{X}_{t}))\|_{F}^{2}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\beta_{t-1}\mathbb{E}\|\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1})\|_{F}^{2}+\frac{\beta_{t-1}^{2}L\eta_{t-1}^{2}n}{1-\beta_{t-1}}+(1-\beta_{t-1})^{2}\sigma^{2}$
$=(1-\alpha_{t})\mathbb{E}\|\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1})\|_{F}^{2}+\alpha_{t}^{2}(2\sqrt{2}Ln+\sigma^{2})$
where ($\circ$) follows from Lemma A.2.
Then, we define $\mathbf{S}_{t}^{\prime}=\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2},\mathbf{S}_{t}=\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2},A_{t+1}=\alpha_{t+1}^{2}(2\sqrt{2}Ln+\sigma^{2})$. Using the conclusion of Lemma A.3, we have
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+A_{t+1}$
Furthermore, since
$\mathbb{E}\|\mathbf{S}_{1}\|$ $=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\mathbf{C}_{1}\|_{F}^{2}=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-(1-\beta_{0})\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}$
$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{0}=0)$
Thus, as in inequality (8), we have
$\sum_{t=1}^{T}\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq 2\sum_{t=1}^{T}(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1})$
$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t+1}$
$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
Then,
$\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}$ $=\|\beta_{t}\mathbf{C}_{t}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}$
$=\|\beta_{t}(\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t}))+(1-\beta_{t})(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t}))\|_{F}$
From this, we can bound the squared norm using the inequality $\|\mathbf{A}+\mathbf{B}\|_{F}^{2}\leq 2\|\mathbf{A}\|_{F}^{2}+2\|\mathbf{B}\|_{F}^{2}$:
$\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\leq 2\beta_{t}^{2}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}^{2}$
Thus, we have
$\mathbb{E}\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}$ $\leq 2\beta_{t}^{2}\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\mathbb{E}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}^{2}$ (11)
$\leq 2\beta_{t}^{2}\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\sigma^{2}$
$\leq 2\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\sigma^{2}$
Thus,
$\sum_{t=1}^{T}\alpha_{t}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}$ $\leq 2\sum_{t=1}^{T}\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\sigma^{2}\sum_{t=1}^{T}\frac{1}{t(t+1)}$ (12)
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}4\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)+2\sigma^{2}$
$=6\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
where $(\circ)$ is due to $\sum_{t=1}^{T}\frac{1}{t(t+1)}=\sum_{t=1}^{T}(\frac{1}{t}-\frac{1}{t+1})=1-\frac{1}{T+1}\leq 1$.
Therefore,
$\Gamma_{t}$ $=\frac{\eta_{t}^{2/3}}{L}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\eta_{t}^{4/3}$
$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}$
Then, we define $A_{1}=8\sqrt{2}Ln+Ln+4L^{-1}\sigma^{2}+L/2$ and $A_{2}=8\sqrt{2}Ln+Ln+10L^{-1}\sigma^{2}+L/2$. We have
$\sum_{t=1}^{T}\Gamma_{t}$ $=\frac{1}{L}\sum_{t=1}^{T}\alpha_{t}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\sum_{t=1}^{T}\alpha_{t}^{2}$ (13)
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{6\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)}{L}+(L/2+Ln)(\ln T+1)$
$\leq A_{1}\ln T+A_{2}$
where ($\circ$) is due to inequality (12).
Then, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $=\frac{1}{T}\sum_{t=1}^{T}t^{3/4}\cdot t^{-3/4}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$
$\leq\frac{1}{T}\sum_{t=1}^{T}t^{3/4}\cdot\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$
$\leq\frac{T^{3/4}}{T}\sum_{t=1}^{T}f(\mathbf{X}_{t})-f(\mathbf{X}_{t+1})+\frac{T^{3/4}}{T}\sum_{t=1}^{T}\Gamma_{t}$
$\leq\frac{f(\mathbf{X}_{1})-f^{*}}{T^{1/4}}+\frac{A_{1}\ln T+A_{2}}{T^{1/4}}$
$=\tilde{\mathcal{O}}(T^{-1/4})$
This completes the proof. ∎
Appendix C Lemmas for Theorem 4.2
C.1 Lemma C.1
Lemma C.1.
Let $\{A_{t}\}_{t\geq 1}$ and $\{B_{t}\}_{t\geq 1}$ be non-negative sequences satisfying the relation $A_{t+1}\leq(1-\varepsilon_{t+1})A_{t}+B_{t+1}$ for all $t\geq 1$. If we define the sequence $\varepsilon_{t}=t^{-p}$ for some constant $p\in(0,1]$, then for all $t\geq 1$, the following inequality holds:
$\sqrt{\varepsilon_{t}}A_{t}\leq 4\left(\frac{A_{t}}{\sqrt{\varepsilon_{t}}}-\frac{A_{t+1}}{\sqrt{\varepsilon_{t+1}}}+\frac{B_{t+1}}{\sqrt{\varepsilon_{t+1}}}\right).$
Proof.
First, we define the function $F(t,q)=\frac{1}{4}t^{-q}-t^{q}+(t+1)^{q}-(t+1)^{-q}$ for $t\geq 1$ and $q\in(0,1/2]$. To analyze its properties, let us define $g(t)=F(t,q)$ for a fixed $q\in(0,1/2]$ and $f(q)=F(t,q)$ for a fixed $t\geq 1$.
Case 1: For $t=1$. We have $g(1)=\frac{1}{4}-1+2^{q}-\frac{1}{2^{q}}$. Since $q\in(0,1/2]$, this expression is bounded above by its value at $q=1/2$, yielding $g(1)\leq\frac{1}{4}-1+2^{1/2}-\frac{1}{2^{1/2}}<0$. Thus, for $t=1$, the inequality $g(t)\leq 0$ holds.
Case 2: For $t>1$. For any given $t>1$ and $q\in(0,1/2]$, we examine the derivative of $f(q)$:
$f^{\prime}(q)$ $=\left(\ln(t+1)(t+1)^{-q}-\frac{1}{4}\ln(t)t^{-q}\right)+(\ln(t+1)(t+1)^{q}-\ln(t)t^{q})$ (14)
$\stackrel{{\scriptstyle(\circ)}}{{\geq}}\left(\ln(t+1)(t+1)^{-q}-\frac{1}{4}\ln(t)t^{-q}\right)$
$=\ln(t)t^{-q}\cdot\left\{\frac{\ln(t+1)}{\ln(t)}\cdot\left(\frac{t}{t+1}\right)^{q}-\frac{1}{4}\right\}$
$\stackrel{{\scriptstyle(\star)}}{{\geq}}\ln(t)t^{-q}\cdot\left\{1\cdot\sqrt{\frac{t}{t+1}}-\frac{1}{4}\right\}$
$\stackrel{{\scriptstyle(\bullet)}}{{\geq}}\ln(t)t^{-q}\cdot\left\{\sqrt{\frac{1}{2}}-\frac{1}{4}\right\}\geq 0$
Here, $(\circ)$ holds because $\ln(t+1)>\ln(t)$ and $(t+1)^{q}>t^{q}$ for all $t\geq 1$ and $q\in(0,1/2]$. The inequality $(\star)$ holds because $\frac{\ln(t+1)}{\ln(t)}>1$ for $t>1$, and the function $(\frac{t}{t+1})^{q}$ is decreasing in $q$, thus its minimum on $(0,1/2]$ is achieved at $q=1/2$. The final inequality $(\bullet)$ holds because $t\geq 1$ implies $\sqrt{\frac{t}{t+1}}\geq\sqrt{\frac{1}{2}}$, and $\sqrt{\frac{1}{2}}>1/4$.
Inequality (14) implies that $f(q)$ is monotonically increasing with respect to $q$ on the interval $(0,1/2]$.
Next, we consider the boundary condition at $q=1/2$. Let $h(t):=f(\frac{1}{2})=\frac{1}{4}t^{-1/2}-t^{1/2}+(t+1)^{1/2}-(t+1)^{-1/2}$. It can be verified that $h(t)\leq 0$ for all $t>1$. Consequently, we have $F(t,\frac{1}{2})=f(\frac{1}{2})\leq 0$ for all $t>1$.
Finally, for all $t>1$ and $q\in(0,1/2]$, we have
$F(t,q)\leq F(t,1/2)\leq 0$ (15)
where the first inequality holds because $f(q)$ is monotonically increasing in $q$ on $(0,1/2]$ for any fixed $t>1$.
With this result, we can proceed as follows:
$\sqrt{\varepsilon_{t}}A_{t}-4\left(\frac{A_{t}}{\sqrt{\varepsilon_{t}}}-\frac{A_{t+1}}{\sqrt{\varepsilon_{t+1}}}+\frac{B_{t+1}}{\sqrt{\varepsilon_{t+1}}}\right)$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\sqrt{\varepsilon_{t}}A_{t}-4\left(\frac{A_{t}}{\sqrt{\varepsilon_{t}}}+\frac{B_{t+1}}{\sqrt{\varepsilon_{t+1}}}\right)+\frac{4}{\sqrt{\varepsilon_{t+1}}}\cdot\left(A_{t}-\varepsilon_{t+1}A_{t}+B_{t+1}\right)$
$=\sqrt{\varepsilon_{t}}A_{t}-4\frac{A_{t}}{\sqrt{\varepsilon_{t}}}+\frac{4}{\sqrt{\varepsilon_{t+1}}}(1-\varepsilon_{t+1})A_{t}$
$=A_{t}\cdot\left(\sqrt{\varepsilon_{t}}-\frac{4}{\sqrt{\varepsilon_{t}}}+\frac{4}{\sqrt{\varepsilon_{t+1}}}-4\sqrt{\varepsilon_{t+1}}\right)$
$\stackrel{{\scriptstyle(\star)}}{{=}}4A_{t}\cdot\left(\frac{1}{4}t^{-q}-t^{q}+(t+1)^{q}-(t+1)^{-q}\right)$
$\stackrel{{\scriptstyle(\bullet)}}{{\leq}}0$
Here, $(\circ)$ follows from the assumption $A_{t+1}\leq(1-\varepsilon_{t+1})A_{t}+B_{t+1}$. The equality $(\star)$ is obtained by substituting $\varepsilon_{t}=t^{-p}$ and setting $q=\frac{p}{2}$ (note that $p\in(0,1]$ implies $q\in(0,1/2]$). Finally, $(\bullet)$ is a direct consequence of our result in inequality (15). ∎
C.2 Lemma C.2
Lemma C.2.
For Algorithm 1 option MVR2, let $\Delta_{t}=\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t};\xi_{t+1})$, $\delta_{t}=\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})$, $\mathbf{S}_{t}=\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})$, and $\mathbf{R}_{t+1}=\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})$. Then we have the following inequality:
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}+P_{t+1}$
where
$A_{t+1}$ $=\frac{B_{t+1}+\beta_{t+1}(\mathbb{E}\|\Delta_{t}\|_{F}^{2}-\|\mathbb{E}\Delta_{t}\|_{F}^{2})}{\mathbb{E}\|\Delta_{t}\|_{F}^{2}}$
$B_{t+1}$ $=(1-\beta_{t+1})\mathbb{E}\langle\Delta_{t},\mathbf{R}_{t+1}\rangle+\beta_{t+1}\mathbb{E}\langle\Delta_{t},\mathbf{S}_{t}\rangle$
$P_{t+1}$ $=\mathbb{E}\|\Delta_{t}\|_{F}^{2}(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1})^{2}-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}$
If we choose $\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{t+1}}$ or $\gamma_{t+1}=1$, then
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$
Proof.
First, we have
$\mathbf{M}_{t+1}$
$=\beta_{t+1}\mathbf{M}_{t}+(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\gamma_{t+1}\cdot\beta_{t+1}(\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t};\xi_{t+1}))$
$=(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\beta_{t+1}(\mathbf{M}_{t}+\gamma_{t+1}(\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t};\xi_{t+1})))$
$=(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\beta_{t+1}(\mathbf{M}_{t}+\gamma_{t+1}\Delta_{t})$
Hence,
$\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})$ $=(1-\beta_{t+1})\left(\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\right)$
$\quad+\beta_{t+1}\left(\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})+\delta_{t}+\gamma_{t+1}\Delta_{t}\right)$
$=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}(\mathbf{S}_{t}+\delta_{t}+\gamma_{t+1}\Delta_{t})$
Note that, according to Assumption 4.4:
$\delta_{t}=\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})=-\mathbb{E}_{t+1}[\Delta_{t}]$
Therefore,
$\mathbf{S}_{t+1}=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\gamma_{t+1}\Delta_{t}-\mathbb{E}\Delta_{t})$
Thus,
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}$ $=\mathbb{E}\|(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\gamma_{t+1}\Delta_{t}-\mathbb{E}\Delta_{t})\|_{F}^{2}$
$=\mathbb{E}\|(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}((\gamma_{t+1}-1)\Delta_{t}+\Delta_{t}-\mathbb{E}\Delta_{t})\|_{F}^{2}$
$=\underbrace{\mathbb{E}\|(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\|_{F}^{2}}_{\text{Term A.1}}$
$\quad+\underbrace{\beta_{t+1}^{2}(\gamma_{t+1}-1)^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}}_{\text{Term A.2}}$
$\quad+\underbrace{2\beta_{t+1}(\gamma_{t+1}-1)\mathbb{E}\langle\Delta_{t},(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\rangle}_{\text{Term A.3}}$
First, let’s consider Term A.1:
A.1 $=\mathbb{E}\|(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\|_{F}^{2}$
$\stackrel{{\scriptstyle(\circ)}}{{=}}\mathbb{E}\|(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
where $(\circ)$ holds because Assumption 4.4 implies $\mathbb{E}[\mathbf{R}_{t+1}]=0$ and $\mathbb{E}[\Delta_{t}-\mathbb{E}\Delta_{t}]=0$, making the cross-terms with $\mathbf{S}_{t}$ zero; and $(\star)$ follows from the inequality $\|\mathbf{A}+\mathbf{B}\|_{F}^{2}\leq 2\|\mathbf{A}\|_{F}^{2}+2\|\mathbf{B}\|_{F}^{2}$ for any $\mathbf{A},\mathbf{B}\in\mathbb{R}^{m\times n}$.
Next, considering the sum of Terms A.2 and A.3, let us define
$A_{t+1}$ $=\frac{B_{t+1}+\beta_{t+1}(\mathbb{E}\|\Delta_{t}\|_{F}^{2}-\|\mathbb{E}\Delta_{t}\|_{F}^{2})}{\mathbb{E}\|\Delta_{t}\|_{F}^{2}}$
$B_{t+1}$ $=(1-\beta_{t+1})\mathbb{E}\langle\Delta_{t},\mathbf{R}_{t+1}\rangle+\beta_{t+1}\mathbb{E}\langle\Delta_{t},\mathbf{S}_{t}\rangle$
Then we have
$\text{A.2}+\text{A.3}$ $=\beta_{t+1}^{2}(\gamma_{t+1}-1)^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}$
$+2(\gamma_{t+1}-1)\beta_{t+1}\mathbb{E}\langle\Delta_{t},(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\rangle$
$=\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}\left((\gamma_{t+1}-1)^{2}+2(\gamma_{t+1}-1)\frac{A_{t+1}}{\beta_{t+1}}\right)$
$=\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}\left(\gamma_{t+1}-1+\frac{A_{t+1}}{\beta_{t+1}}\right)^{2}-\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}\left(\frac{A_{t+1}}{\beta_{t+1}}\right)^{2}$
$=\mathbb{E}\|\Delta_{t}\|_{F}^{2}(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1})^{2}-A_{t+1}^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}$
$=P_{t+1}$
Therefore, combining the bounds:
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}$ $\leq 2(1-\beta_{t+1})^{2}\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}2(1-\beta_{t+1})^{2}\sigma^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\sigma^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1}$
where $(\circ)$ uses Assumption 4.4 (bounded variance, i.e., $\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}\leq\sigma^{2}$) and the property that $\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}\leq\mathbb{E}\|\Delta_{t}\|_{F}^{2}$; and $(\star)$ follows from Assumption 4.3, which implies $\|\Delta_{t}\|_{F}^{2}\leq L^{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$.
Next, we set
$\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{t+1}}\text{ or }\gamma_{t+1}=\gamma=1$ (16)
Then,
$P_{t+1}=-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}\leq 0\text{ or }P_{t+1}=0$
This leads to the final result:
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}$ $\leq\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$
where $(\circ)$ follows from $\beta_{t+1}^{2}\leq\beta_{t+1}\leq 1$. ∎
C.3 Lemma C.3
Lemma C.3.
Let $\{\mathbf{S}_{t}\}_{t\geq 1}$ be a sequence of matrices satisfying the recursive inequality $\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})$ for some constants $L,n,\sigma^{2}>0$. If we set the step size $\eta_{t}=t^{-2/3}$ and $\gamma=1$, then we have the following upper bound on the time-averaged expectation:
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}.$
Proof.
The proof begins with the recursive inequality derived from a preceding Lemma C.2:
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t}^{2}L^{2}n+2\eta_{t+1}^{2}\sigma^{2}$
where $\mathbf{S}_{t+1}=\nabla f(\mathbf{X}_{t+1})-\mathbf{M}_{t+1}$. We have noticed the following facts
$\frac{1}{t^{2/3}}\leq\frac{2}{(t+1)^{2/3}}$
Therefore , we have
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})$
Let $A_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ and $B_{t+1}=2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})$. The inequality can be written as $A_{t+1}\leq(1-\eta_{t+1})A_{t}+B_{t+1}$. This structure allows us to apply a standard result Lemma C.1 which yields:
$\sqrt{\eta_{t}}A_{t}\leq 4\left(\frac{A_{t}}{\sqrt{\eta_{t}}}-\frac{A_{t+1}}{\sqrt{\eta_{t+1}}}+\frac{B_{t+1}}{\sqrt{\eta_{t+1}}}\right)$
Then, we define $P_{t}=\frac{4A_{t}}{\sqrt{\eta_{t}}}=\frac{4\mathbb{E}\|\mathbf{S}_{t}\|^{2}_{F}}{\sqrt{\eta_{t}}}$. Substituting $P_{t}$ and the definition of $B_{t+1}$ into the inequality gives:
$\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq P_{t}-P_{t+1}+\frac{4\cdot 2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})}{\sqrt{\eta_{t+1}}}=P_{t}-P_{t+1}+8\eta_{t+1}^{3/2}(2L^{2}n+\sigma^{2})$
Now, we sum this inequality from $t=1$ to $T$:
$\sum_{t=1}^{T}\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq\sum_{t=1}^{T}\left(P_{t}-P_{t+1}+8\eta_{t+1}^{3/2}(2L^{2}n+\sigma^{2})\right)$
$=(P_{1}-P_{T+1})+8(2L^{2}n+\sigma^{2})\sum_{t=1}^{T}\eta_{t+1}^{3/2}$
Since $P_{T+1}\geq 0$, we can drop this term to simplify the bound. By setting the step size $\eta_{t}=t^{-2/3}$, we have $\eta_{t+1}^{3/2}=((t+1)^{-2/3})^{3/2}=(t+1)^{-1}$. The summation becomes:
$\sum_{t=1}^{T}\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq P_{1}+8(2L^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t+1}$ (17)
$\leq P_{1}+8(2L^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)$
where $(\circ)$ follows from the harmonic series, $\sum_{t=1}^{T}\frac{1}{t}\leq 1+\ln T$.
Finally, we establish the bound on the time-averaged expectation. With our choice of $\eta_{t}=t^{-2/3}$, we have $\sqrt{\eta_{t}}=t^{-1/3}$. Therefore:
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $=\frac{1}{T}\sum_{t=1}^{T}t^{1/3}\cdot t^{-1/3}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
$=\frac{1}{T}\sum_{t=1}^{T}t^{1/3}\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
$\leq\frac{T^{1/3}}{T}\sum_{t=1}^{T}\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\text{(since }t^{1/3}\leq T^{1/3}\text{ for }t\leq T\text{)}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{T^{1/3}}{T}\left(P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)\right)$
$=\frac{P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}$
where $(\circ)$ follows from the inequality (17).
Substituting $P_{1}=4\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}/\sqrt{\eta_{1}}=4\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}$ completes the main proof. To establish the final bound, we now analyze the initial term $P_{1}$. Given the definition $\mathbf{S}_{1}=\nabla f(\mathbf{X}_{1})-\mathbf{M}_{1}$, we have:
$P_{1}$ $=4\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\mathbf{M}_{1}\|_{F}^{2}$
$=4\mathbb{E}\|\nabla f(\mathbf{X}_{1})-(1-\beta_{1}+\gamma_{1}\beta_{1})\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}$
$=4\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}4\sigma^{2}$
The final inequality $(\star)$ holds by setting the parameter $\gamma_{1}=\gamma=1$. This choice nullifies the first term, as $(\gamma_{1}-1)^{2}=0$, and simplifies the coefficient of the variance to $(1-\beta_{1}+\beta_{1})^{2}=1$. ∎
Appendix D Proofs of Theorem 4.2
Proof.
According to Assumption 4.3, and based on the descent lemma, we have
$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})+\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{t})+\langle\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{*}+\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{*}+\frac{1}{2\alpha}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\alpha+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
where $(\circ)$ holds because, by the definition of $\mathbf{O}_{t}$ and the property of norms $\|\mathbf{M}_{t}\|_{*}\geq\|\mathbf{M}_{t}\|_{F}$, we have:
$\langle\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}$
and $(\star)$ holds by setting $\alpha=\frac{L}{\sqrt{\eta_{t}}}$. Thus, we have
$\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$ $\leq\sum_{t=1}^{T}\left(\mathbb{E}[f(\mathbf{X}_{t})]-\mathbb{E}[f(\mathbf{X}_{t+1})]\right)+\sum_{t=1}^{T}\frac{\sqrt{\eta_{t}}}{2L}\mathbb{E}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}$
$\quad+\sum_{t=1}^{T}\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{1})-f^{*}+\sum_{t=1}^{T}\frac{\sqrt{\eta_{t}}}{2L}\mathbb{E}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\sum_{t=1}^{T}\frac{L(\eta_{t}^{3/2}+\eta_{t}^{2})}{2}n$
$\leq f(\mathbf{X}_{1})-f^{*}+\frac{1}{2L}\sum_{t=1}^{T}\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{Ln}{2}\sum_{t=1}^{T}t^{-1}+\frac{Ln}{2}\sum_{t=1}^{T}t^{-4/3}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}f(\mathbf{X}_{1})-f^{*}+\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{2L}$
$\quad+\frac{Ln}{2}(1+\ln T)+\frac{Ln}{2}\sum_{t=1}^{T}t^{-4/3}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+2Ln+4(2Ln+\sigma^{2}L^{-1})(1+\ln T)+\frac{Ln}{2}(1+\ln T)$
where $(\circ)$ uses Lemma C.2; $(\star)$ follows from the fact that $\sum_{t=1}^{T}\frac{1}{t^{4/3}}\leq 4$. Next, we let
$G=f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+\left(8Ln+4\sigma^{2}L^{-1}+\frac{Ln}{2}\right)(1+\ln T)+2Ln$
Thus, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\frac{t^{2/3}}{t^{2/3}}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$
$\leq\frac{T^{2/3}}{T}\sum_{t=1}^{T}\frac{1}{t^{2/3}}\mathbb{E}\|\mathbf{M}_{t}\|_{F}=\frac{1}{T^{1/3}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$
$\leq\frac{G}{T^{1/3}}$
Next, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}$ $\stackrel{{\scriptstyle(\circ)}}{{\leq}}\sqrt{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}\sqrt{\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}}$
where $(\circ)$ uses Jensen’s inequality; $(\star)$ uses Lemma C.3 by letting $\mathbf{S}_{t}=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}$. Thus, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$
$\leq\sqrt{\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}}+\frac{G}{T^{1/3}}$
$=\mathcal{O}\left(\frac{\ln T}{T^{1/3}}\right)$
This completes the proof. ∎
Appendix E Lemmas for Theorems 4.3 and 4.4
E.1 Lemma E.1
Lemma E.1.
Let $\{\Gamma_{t}\}_{t\geq 1}$ be a non-negative sequence whose partial sum $D_{t}=\sum_{t=1}^{T}\Gamma_{t}$ is bounded by $D_{t}\leq A_{1}\ln T+A_{2}$ for some positive constants $A_{1}$ and $A_{2}$. Define the function $f(t)=\frac{t^{1-p}}{\ln t}$ where $p\in(0,1)$. Then, for any sufficiently large integer $T_{0}>e^{1/(1-p)}$ where $e$ is the base of the natural logarithm, the following inequality holds for all $T>T_{0}$:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq\left(A_{1}+A_{2}\right)(T-1)^{1-p}$
Proof.
We choose an integer $T_{0}>e^{1/(1-p)}$ large enough such that for all $t\geq T_{0}$, the function $f(t)$ is positive and monotonically increasing. We use summation by parts to bound the sum $\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)$:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)=\sum_{t=T_{0}}^{T-1}(D_{t}-D_{t-1})f(t)=D_{T-1}f(T-1)-D_{T_{0}-1}f(T_{0})-\sum_{t=T_{0}}^{T-1}D_{t-1}(f(t)-f(t-1))$
Since $D_{t-1}\geq 0$ and $f(t)$ is monotonically increasing for $t\geq T_{0}$ (i.e., $f(t)-f(t-1)\geq 0$), the last term, $\sum_{t=T_{0}}^{T-1}D_{t-1}(f(t)-f(t-1))$, is non-negative. Therefore, we have:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq D_{T-1}f(T-1)$
Substituting the given bound on the partial sums, $D_{T-1}\leq A_{1}\ln(T-1)+A_{2}$, we obtain:
$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)$ $\leq(A_{1}\ln(T-1)+A_{2})\frac{(T-1)^{1-p}}{\ln(T-1)}$
$=\left(A_{1}+\frac{A_{2}}{\ln(T-1)}\right)(T-1)^{1-p}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}(A_{1}+A_{2})(T-1)^{1-p}$
where $(\circ)$ follows from $\ln(T-1)\geq 1$ for sufficiently large $T$.
Consequently, for any given $\epsilon>0$, the expression can be bounded by $(A_{1}+A_{2}+\epsilon)T^{1-p}$ for sufficiently large $T$. ∎
E.2 Lemma E.2
Lemma E.2.
Suppose a positive sequence $\{\Delta_{t}\}_{t\geq 1}$ satisfies the following recursive inequality:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}$
where $p\in(0,1)$ and $\mu>0$ are constants. The non-negative noise sequence $\{\Gamma_{t}\}_{t\geq 1}$ satisfies the condition from Lemma E.1, with its partial sum bounded by $\sum_{t=1}^{T}\Gamma_{t}=\mathcal{O}(\ln T)$.
Then, the sequence $\Delta_{t}$ converges to 0 with the following rate:
$\Delta_{t}=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right)$
Proof.
We first define the variable $G_{t}=\sqrt{\Delta_{t}}$. Since $\Delta_{t}>0$, it follows that $G_{t}>0$. The function $f(x)=\sqrt{x}$ is concave, which implies $\sqrt{y}-\sqrt{x}\leq\frac{1}{2\sqrt{x}}(y-x)$ for any $x,y>0$. By setting $y=\Delta_{t+1}$ and $x=\Delta_{t}$, we obtain:
$G_{t+1}-G_{t}\leq\frac{\Delta_{t+1}-\Delta_{t}}{2G_{t}}$
Substituting the original recurrence relation $\Delta_{t+1}-\Delta_{t}\leq-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}$ and noting that $\sqrt{\Delta_{t}}=G_{t}$, we have:
$G_{t+1}-G_{t}\leq\frac{1}{2G_{t}}\left(-\frac{\sqrt{2\mu}}{t^{p}}G_{t}+\Gamma_{t}\right)$
Letting $C=\frac{\sqrt{2\mu}}{2}$, we arrive at the core recurrence for $G_{t}$:
$G_{t+1}\leq G_{t}-\frac{C}{t^{p}}+\frac{\Gamma_{t}}{2G_{t}}$ (18)
For $p\in(0,1)$, we will prove that $G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1-p}}\right)$ by contradiction.
Assume the proposition is false. This means that for any constant $A>0$, there exist infinitely many time steps $t$ such that $G_{t}>B_{t}$, where $B_{t}=A\cdot\frac{\ln t}{t^{1-p}}$.
We choose a constant $A$ large enough (the specific condition will be derived later). By our assumption, there must exist an arbitrarily large time $T$ such that $G_{T}>B_{T}$. Consider an interval $T_{0},T-1]$ and assume that $G_{t}>B_{t}$ for all $t$ in this interval.
Summing the inequality (18) from $t=T_{0}$ to T1T-1 gives:
$G_{T}-G_{T_{0}}\leq\sum_{t=T_{0}}^{T-1}(G_{t+1}-G_{t})\leq\sum_{t=T_{0}}^{T-1}\left(-\frac{C}{t^{p}}+\frac{\Gamma_{t}}{2G_{t}}\right)$
Over this interval, since $G_{t}>B_{t}=A\frac{\ln t}{t^{1-p}}$, we have $\frac{1}{G_{t}}<\frac{1}{B_{t}}=\frac{t^{1-p}}{A\ln t}$. Substituting this bound yields:
$G_{T}<G_{T_{0}}-\underbrace{C\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}}_{V_{1}}+\underbrace{\frac{1}{2A}\sum_{t=T_{0}}^{T-1}\frac{\Gamma_{t}t^{1-p}}{\ln t}}_{V_{2}}$
Next, we analyze the asymptotic behavior of the two sums, $V_{1}$ and $V_{2}$.
For the negative drift term $V_{1}$, we use an integral approximation:
$\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}\geq\int_{T_{0}}^{T}x^{-p}dx=\frac{1}{1-p}(T^{1-p}-T_{0}^{1-p})$
The leading term of $V_{1}$ is thus on the order of $\frac{C}{1-p}T^{1-p}$.
For the positive noise term $V_{2}$, we use the condition on the partial sums of $\Gamma_{t}$, $S_{t}=\sum_{k=1}^{t}\Gamma_{k}\leq A_{1}\ln t+A_{2}$. Applying summation by parts (in a manner analogous to Lemma E.1), the dominant behavior of the sum in $V_{2}$ is given by $S_{T-1}\cdot\frac{(T-1)^{1-p}}{\ln(T-1)}$. This can be bounded by $((A_{1}+A_{2})\ln(T-1))\frac{(T-1)^{1-p}}{\ln(T-1)}=(A_{1}+A_{2})(T-1)^{1-p}$. For simplicity in asymptotics, let us denote the effective constant as $\mathcal{A}:=A_{1}+A_{2}$. The leading behavior of $V_{2}$ is thus on the order of $\frac{\mathcal{A}}{2A}T^{1-p}$.
Substituting the leading terms of these bounds back into the inequality for $G_{T}$:
$G_{T}<G_{T_{0}}-\frac{C}{1-p}T^{1-p}+\frac{\mathcal{A}}{2A}T^{1-p}=G_{T_{0}}+T^{1-p}\left(\frac{\mathcal{A}}{2A}-\frac{C}{1-p}\right)$
We now choose the constant $A$ such that the coefficient of the leading term is negative:
$\frac{\mathcal{A}}{2A}-\frac{C}{1-p}<0\implies A>\frac{\mathcal{A}}{2C}$
With an $A$ satisfying this condition, for a sufficiently large $T$, the negative term proportional to $T^{1-p}$ will dominate, forcing the right-hand side to become negative. This contradicts the fact that $G_{T}=\sqrt{\Delta_{T}}$ must be positive.
This contradiction proves that our assumption that $G_{t}$ can remain above $B_{t}$ for an arbitrarily long interval is false. Therefore, there must exist a time $T_{A}$ such that for all $t>T_{A}>T_{0}$, $G_{t}\leq B_{t}$. This establishes that $G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1-p}}\right)$.
Finally, squaring both sides of this result yields the convergence rate for $\Delta_{t}$:
$\Delta_{t}=G_{t}^{2}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1-p}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right)$
Appendix F Proofs of Theorem 4.3
Proof.
By Theorem 4.1 inequality (7), we have
$\mathbb{E}[f(\mathbf{X}_{t+1})]\leq\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]+\Gamma_{t}$ (19)
where
$\Gamma_{t}=\frac{\eta_{t}^{2/3}}{L}\mathbb{E}\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+(L/2+Ln)\eta_{t}^{4/3}.$
Case 1: $\gamma=0$.
Let $\eta_{t}=t^{-3/4},\beta_{t}=1-t^{-1/2},\alpha_{t}=t^{-1/2}$. According to Theorem 4.1 inequality (9), we have
$\sum_{t=1}^{T}\Gamma_{t}\leq A_{1}\ln T+A_{2}$
where $A_{1}=2L^{-1}\sigma^{2}+4\sqrt{2}Ln+Ln+L/2$ and $A_{2}=4L^{-1}\sigma^{2}+4\sqrt{2}Ln+Ln+L/2$.
Case 2: $\gamma\neq 0$.
Let $\eta_{t}=t^{-3/4},\beta_{t}=1-(1+t)^{-1/2},\alpha_{t}=t^{-1/2}$. According to Theorem 4.1 inequality (13), we have
$\sum_{t=1}^{T}\Gamma_{t}\leq A_{1}\ln T+A_{2}$
where $A_{1}=4L^{-1}\sigma^{2}+8\sqrt{2}Ln+Ln+L/2$ and $A_{2}=10L^{-1}\sigma^{2}+8\sqrt{2}Ln+Ln+L/2$.
Then, we define $\Delta_{t}=\mathbb{E}[f(\mathbf{X}_{t})]-f^{*}$ . Applying Assumption 4.5 to inequality (19), we have
$\Delta_{t+1}$ $\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\Delta_{t}}+\Gamma_{t}$
$=\Delta_{t}-\frac{\sqrt{2\mu}}{t^{3/4}}\sqrt{\Delta_{t}}+\Gamma_{t}$
By Lemma E.2, we have
$G_{t+1}\leq G_{t}-\frac{C}{t^{3/4}}+\frac{\Gamma_{t}}{2G_{t}}$
Setting $p=3/4$, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{8C}$, there exists a time step $T_{A}$ such that for all $t>T_{A}$, we have $G_{t}\leq A\cdot\frac{\ln t}{t^{1/4}}$. Thus, we obtain the convergence rate for $G_{t}$:
$G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1/4}}\right)$
Finally, since $\Delta_{t}=G_{t}^{2}$, squaring both sides yields the final convergence rate of $\Delta_{t}$:
$\Delta_{t}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1/4}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{1/2}}\right)$
Hence,
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right)$
This completes the proof. ∎
Appendix G Proofs of Theorem 4.4
Proof.
According to Theorem 4.2, we have
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
Since $\eta_{t}\leq 1$, we have $\frac{L}{\sqrt{\eta_{t}}}\geq L$. Thus, taking the expectation yields
$\mathbb{E}[f(\mathbf{X}_{t+1})]$ $\leq\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\mathbf{M}_{t}\|_{F}]+\frac{\sqrt{\eta_{t}}}{2L}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+\frac{L}{\sqrt{\eta_{t}}}\mathbb{E}[\|\mathbf{X}_{T+1}-\mathbf{X}_{t}\|_{F}^{2}]$
$\leq\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\mathbf{M}_{t}\|_{F}]+\frac{\sqrt{\eta_{t}}}{2L}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]+\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}]$
$+\frac{\sqrt{\eta_{t}}}{2L}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}$
$\stackrel{{\scriptstyle(\star)}}{{\leq}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]+\frac{1}{2\epsilon}\eta_{t}^{2}+\frac{\epsilon}{2}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]$
$+\frac{\sqrt{\eta_{t}}}{2L}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}$
$\stackrel{{\scriptstyle(\bullet)}}{{=}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]$
$+\underbrace{\frac{L\eta_{t}^{3/2}}{2}+\left(\frac{\epsilon}{2}+\frac{\sqrt{\eta_{t}}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}}_{\Gamma_{t}}$
where $(\circ)$ follows from the reverse triangle inequality $\|\mathbf{M}_{t}\|_{F}\leq\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}-\|\nabla f(\mathbf{X}_{t})\|_{F}$; $(\star)$ applies Young’s inequality to the term $\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}]$; and $(\bullet)$ collects the residual terms into $\Gamma_{t}$. By setting $\epsilon=\frac{\sqrt{\eta_{t}}}{L}$, we can simplify the expression for $\Gamma_{t}$:
$\Gamma_{t}=\frac{\sqrt{\eta_{t}}}{L}\mathbb{E}[\|\mathbf{S}_{t}\|_{F}^{2}]+(L/2+Ln)\eta_{t}^{3/2}$
where $\mathbf{S}_{t}=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}$. The sum of these terms can be bounded. Let $A_{1}=17Ln+8L^{-1}\sigma^{2}+L/2$ and $A_{2}=17Ln+12L^{-1}\sigma^{2}+L/2$. We have
$\sum_{t=1}^{T}\Gamma_{t}$ $=\frac{1}{L}\sum_{t=1}^{T}\sqrt{\eta_{t}}\mathbb{E}[\|\mathbf{S}_{t}\|_{F}^{2}]+\sum_{t=1}^{T}(L/2+Ln)\eta_{t}^{3/2}$
$\leq\frac{1}{L}\sum_{t=1}^{T}\sqrt{\eta_{t}}\mathbb{E}[\|\mathbf{S}_{t}\|_{F}^{2}]+(L/2+Ln)\sum_{t=1}^{T}\frac{1}{t}$
$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{L}+(L/2+Ln)(1+\ln T)$
$=A_{1}\ln T+A_{2}$
where $(\circ)$ holds by Lemma C.3 inequality (17).
Next, let $\Delta_{t}=\mathbb{E}[f(\mathbf{X}_{t})]-f^{*}$. By applying the Assumption 4.5, $\|\nabla f(\mathbf{X}_{t})\|_{F}^{2}\geq 2\mu(f(\mathbf{X}_{t})-f^{*})$, and Jensen’s inequality to the main recurrence, we obtain
$\Delta_{t+1}$ $\leq\Delta_{t}-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]+\Gamma_{t}$
$\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\mathbb{E}[f(\mathbf{X}_{t})-f^{*}]}+\Gamma_{t}$
$\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\Delta_{t}}+\Gamma_{t}$
Setting the step size $\eta_{t}=1/t^{2/3}$, the recurrence becomes
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{2/3}}\sqrt{\Delta_{t}}+\Gamma_{t}$
This recurrence is in the form required by Lemma E.2. Let $G_{t}=\sqrt{\Delta_{t}}$, $C=\sqrt{2\mu}/2$ and $p=2/3$. According to Lemma E.2, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{6C}$, there exists a time $T_{A}$ such that for all $t>T_{A}$, $G_{t}\leq A\cdot\frac{\ln t}{t^{1/3}}$. This establishes the convergence rate for $G_{t}$:
$G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1/3}}\right)$
Finally, since $\Delta_{t}=G_{t}^{2}$, squaring both sides gives the final convergence rate for $\Delta_{t}$:
$\Delta_{t}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1/3}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2/3}}\right)$
Therefore, we have
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right)$
This completes the proof. ∎
Appendix H Experimental Details
H.1 Training on CIFAR10
The ResNet18 model He et al. (2016) undergoes pretraining on the CIFAR-10 dataset with comprehensive hyperparameter specifications provided in Table 2. For each optimizer, the learning rate is selected via a grid search over the set $\{1\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}\}$. To ensure a robust comparison, all experiments are repeated over five different random seeds, and we report the mean results with one standard deviation shaded.
Table 2: Hyperparameters used for training ResNet18 on CIFAR10
SGD Adam Muon-Type
Model Size 42.7M
Training Epochs 100
Batch Size 128
Learning Rate 0.1 0.01 0.05
Learning Rate Scheduling cosine to 10%
Numerical precision float32
Weight Decay 0.01
$(\beta_{1},\beta_{2})$ (0.9,0.999) (0.9,-)
Gamma 0.1
H.2 Pretraining on C4
The LLaMA2-130M modelTouvron et al. (2023) was pre-trained on the C4 (Colossal Clean Crawled Corpus) dataset Raffel et al. (2020). Identical learning rates and scheduling protocols were systematically implemented across all optimizers during the training process. Comprehensive experimental specifications are tabulated in Table 3. We use the Muon implementation from Moonlight [^1] https://github.com/MoonshotAI/Moonlight.
Table 3: Hyperparameters used for training LLaMA2-130M on C4
AdamW Muon-Type
Model Size 130M
Hidden Size 768
Head 12
Depth 12
Training Steps 10000
Warmup Steps 1000
Maximum Length 1024
Batch Size 512
Learning Rate 1e-3
Warmup Scheduling linear from 8e-6
Learning Rate Scheduling cosine to 10%
Numerical precision bfloat16
Weight Decay 0.1
$(\beta_{1},\beta_{2})$ (0.9,0.999) (0.95,-)
Gamma 0.05

### v3

@@
 The Muon optimizer has demonstrated remarkable empirical success in handling matrix-structured parameters for training neural networks.
-However, a significant gap persists between its practical performance and theoretical understanding.
-Existing analyses indicate that the standard Muon variant achieves only a suboptimal convergence rate of $\mathcal{O}(T^{-1/4})$ in stochastic non-convex settings, where $T$ denotes the number of iterations.
-To explore the theoretical limits of the Muon framework, we develop and analyze two momentum-based variance-reduced variants: a one-batch version (Muon-MVR1) and a two-batch version (Muon-MVR2).
-We provide the first rigorous proof that incorporating a variance-reduction mechanism enables Muon-MVR2 to attain an optimal convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$, thereby matching the theoretical lower bound for this class of problems.
-Moreover, our analysis establishes convergence guarantees for Muon variants under the Polyak-Łojasiewicz (PŁ) condition.
+However, a significant gap remains between its practical performance and theoretical understanding.
+Existing analyses show that the Muon variants achieve only a suboptimal iteration complexity of $\mathcal{O}(T^{-1/4})$ in stochastic non-convex settings, where $T$ denotes the number of iterations.
+To explore the theoretical limits of the Muon framework, we analyze two Momentum-based Variance-Reduced variants: a one-batch version (Muon-MVR1) and a two-batch version (Muon-MVR2).
+We provide the first rigorous proof that incorporating variance reduction enables Muon-MVR2 to attain the optimal iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$, thereby matching the theoretical lower bound for this class of problems.
+Furthermore, our analysis establishes last-iterate convergence guarantees for Muon variants under the Polyak-Łojasiewicz (PŁ) condition.
 Extensive experiments on vision (CIFAR-10) and language (C4) benchmarks corroborate our theoretical findings on per-iteration convergence.
-Overall, this work provides the first proof of optimality for a Muon-style optimizer and clarifies the path toward developing more practically efficient, accelerated variants.
+Overall, this work offers the first proof of optimality for a Muon-style optimizer and clarifies the path toward developing more practically efficient, accelerated variants.
 1 Introduction
-The immense computational cost of pre-training Large Language Models (LLMs) has spurred a surge of research into novel optimization methods designed to enhance parameter efficiency and training stability Hoffmann et al.
-(2022); Liu et al.
-(2023); Chen et al.
-(2023); Vyas et al.
-(2025); Pethick et al.
-(2025); Yuan et al.
-(2024).
-Among these, methods based on matrix orthogonalization have recently garnered significant attention from both academia and industry Jordan et al.
-(2024); Liu et al.
-(2025a).
-In particular, the Muon optimizer has emerged as a notable milestone due to its impressive empirical performance Liu et al.
-(2025a); An et al.
-(2025); Liu et al.
-(2025b); Shah et al.
-(2025).
+The immense computational cost of pre-training Large Language Models (LLMs) has spurred a surge of research into novel optimization methods designed to enhance parameter efficiency and training stability hoffmann2022training; Liu2023SophiaAS; Chen2023SymbolicDO; vyas2025soap; pethick2025training; Yuan2024MARSUT.
+Among these, methods based on matrix orthogonalization have recently garnered significant attention from both academia and industry jordan2024muon; Liu2025MuonIS.
+In particular, the Muon optimizer has emerged as a notable milestone due to its impressive empirical performance Liu2025MuonIS; An2025ASGOAS; liu2025cosmos; shah2025practical.
 However, despite its practical success, the theoretical understanding of Muon’s underlying mechanisms has surprisingly lagged behind, with existing convergence analyses being fraught with limitations and even critical fallacies.
 Specifically, the current theoretical exploration of Muon’s convergence faces three primary obstacles.
-First, some analyses are predicated on flawed mathematical assumptions, such as the erroneous use of the inequality $\|\mathbf{S}^{-1}\|_{2}\leq 1/\|\mathbf{S}\|_{2}$ in a key part of their proof, which casts serious doubt on the validity of their convergence claims Li & Hong (2025).
-Second, existing analyses have not established convergence to a stationary point without reliance on problem dimension or batch size; the available bounds become valid only when the batch size is sufficiently large, so the basic requirement for a trustworthy optimizer remains unfulfilled Sato et al.
-(2025); Sfyraki & Wang (2025).
+First, existing analyses have not established convergence to a stationary point without reliance on problem dimension or batch size; the available bounds become valid only when the batch size is sufficiently large, so the basic requirement for a trustworthy optimizer remains unfulfilled Sato2025ConvergenceBA; sfyraki2025lions.
+Second, some analyses are predicated on flawed mathematical assumptions, such as the erroneous use of the inequality $\|\mathbf{S}^{-1}\|_{2}\leq 1/\|\mathbf{S}\|_{2}$ in a key part of their proof, which casts serious doubt on the validity of their convergence claims Li2025ANO.
 Finally, the most rigorous existing work analyzes both the standard Muon algorithm (like Algorithm 1, Option MVR1 with $\gamma=0$) and its simplified variant , considering their convergence to a non-standard $\epsilon$-nuclear norm stationary point.
-This leaves the behavior of the Nesterov-Accelerated Muon algorithm (like Algorithm 1, Option MVR1), as discussed in Liu et al.
-(2025a); Sato et al.
-(2025), and Variance-Reduction Muon (like Algorithm 1, Option MVR2) under standard settings as an open question Shen et al.
-(2025).
+This leaves the behavior of the Nesterov-Accelerated Muon algorithm (like Algorithm 1, Option MVR1), as discussed in Liu2025MuonIS; Sato2025ConvergenceBA, and Variance-Reduction Muon (like Algorithm 1, Option MVR2) under standard settings as an open question shen2025convergence.
 To bridge this critical theoretical gap, this work establishes a rigorous and complete theoretical foundation for the Muon optimizer. Our main contributions are threefold:
   • For general non-convex settings, we provide a convergence analysis for both the standard Muon (Algorithm 1, Option MVR1 with $\gamma=0$) and the Muon-MVR1 algorithm (Algorithm 1, Option MVR1).
-  • Regarding the optimal convergence rate, we are the first to demonstrate that the Variance-Reduction Muon-MVR2 algorithm (Algorithm 1, Option MVR2) can overcome the bottleneck of traditional momentum methods, achieving the optimal ergodic convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$.
+  • Regarding the optimal iteration complexity, we are the first to demonstrate that the Variance-Reduction Muon-MVR2 algorithm (Algorithm 1, Option MVR2) can overcome the bottleneck of traditional momentum methods, achieving the optimal iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$.
   • Under the Polyak-Łojasiewicz (PL) condition, we prove that our proposed algorithms, Muon-MVR1, and Muon-MVR2, all achieve sublinear convergence rates.
-More precisely, we demonstrate that Muon-MVR1 converge at a rate of $\tilde{\mathcal{O}}(1/T^{1/2})$, and Muon-MVR2 converges at an accelerated rate of $\tilde{\mathcal{O}}(1/T^{2/3})$
-We also refer readers to see the main contributions of our works and comparisons with the existing works in Table 1.
-Table 1: Comparison for existing Muon-type analyses with ours
-Smooth
+More precisely, we demonstrate that Muon-MVR1 converge at a rate of $\tilde{\mathcal{O}}(T^{-1/2})$, and Muon-MVR2 converges at an accelerated rate of $\tilde{\mathcal{O}}(T^{-2/3})$
+Table 1 summarizes the main contributions of our work and compares them with existing methods.
+Smootha
 Stoc. Gradient
-Estimator
+Estimator b
 Batch
@@
 Complexity
-Convergence
-Rate
-Li & Hong (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
-Sato et al. (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(n)$
-Sato et al. (2025) $L$ Option MVR1 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(n)$
-Sfyraki & Wang (2025) $L_{+}$ Option MVR2 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(T^{-1/3})$
-Shen et al. (2025) $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
-Ours $L$ Option MVR1($\gamma=0$) $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ Th4.3
-Ours $L$ Option MVR1 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ Th4.3
-Ours $L_{+}$ Option MVR2 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/3})$ Th4.4
+Last-iterate
+Conv. Rate
+Li2025ANO $L$ MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
+Sato2025ConvergenceBA $L$ MVR1($\gamma=0$) $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(T^{-1})+\mathcal{O}(1)^{c}$
+Sato2025ConvergenceBA $L$ MVR1 $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(T^{-1})+\mathcal{O}(1)^{c}$
+sfyraki2025lions $L_{+}$ MVR2 $\mathcal{O}(\epsilon^{-1})^{d}$ $\mathcal{O}(T^{-1/3})$
+shen2025convergence $L$ MVR1($\gamma=0$) $\mathcal{O}(1)$ $\mathcal{O}(T^{-1/4})$
+Ours $L$ MVR1($\gamma=0$) $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ $\tilde{\mathcal{O}}(T^{-1/2})$
+Ours $L$ MVR1 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/4})$ $\tilde{\mathcal{O}}(T^{-1/2})$
+Ours $L_{+}$ MVR2 $\mathcal{O}(1)$ $\tilde{\mathcal{O}}(T^{-1/3})$ $\tilde{\mathcal{O}}(T^{-2/3})$
   • a
+    $L$ pertains to the smoothness of the overall function $f$, while $L_{+}$ pertains to the smoothness of its stochastic components $f(\cdot;\xi)$.
+  • b
     Option MVR1 ($\gamma=0$) is the standard momentum method, Option MVR1 is the one-batch variance-reduction momentum method, and Option MVR2 is the two-batch variance-reduction momentum method.
 We summarize them in Algorithm 1.
-  • b
-    Although increasing the batch size can mitigate the impact of stochastic noise Sato et al.
-(2025), these methods still fail to converge to a stationary point and cannot eliminate the influence of dimensionality.
   • c
-    The results from Sfyraki & Wang (2025) on Option MVR2 are the closest to ours.
+    Although increasing the batch size can mitigate the impact of stochastic noise Sato2025ConvergenceBA, these methods still fail to converge to a stationary point and cannot eliminate the influence of dimensionality.
+  • d
+    The results from sfyraki2025lions on Option MVR2 are the closest to ours.
 However, their method requires a large initial batch of size $\mathcal{O}(\epsilon^{-1})$, although the batch size can be reduced to 1 in subsequent iterations.
-    $L$ pertains to the smoothness of the overall function $f$, while $L_{+}$ pertains to the smoothness of its stochastic components $f(\cdot;\xi)$.
-2 Preliminaries and Problem Setup
-2.1 Notations
+Table 1: Comparison of Existing Muon-type Analyses with Ours.
+$\bullet$ Organizations.
+The rest of the paper is organized as follows.
+Section 2 reviews existing variants of the Muon algorithm.
+Section 3 addresses iteration complexity and last-iterate convergence rate.
+Section 4 presents experimental results demonstrating the effectiveness of our method.
+Related work is reviewed in Section 5, and conclusions are drawn in Section 6.
+$\bullet$ Notations.
 We denote scalars by non-bold letters (e.g., $a,A$), vectors in $\mathbb{R}^{d}$ by bold lowercase letters (e.g., $\mathbf{a}$), and matrices by bold uppercase letters (e.g., $\mathbf{A}$).
@@
 Let $\xi$ be a random variable that is independent of $\mathbf{X}\in\mathbb{R}^{m\times n}$, and let $\mathbb{E}_{\xi}[\mathbf{X}]$ denote the expectation of a random variable $\mathbf{X}$ with respect to $\xi$.
-2.2 Problem Setup
-We consider minimizing the objective function as follows
-$\min_{\mathbf{X}}f(\mathbf{X}),\mathrm{~where~}f(\mathbf{X})=\mathbb{E}_{\xi\sim\mathcal{D}}[f(\mathbf{X};\xi)].$ (1)
-For problem (1), let $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ be a function where $\mathbf{X}\in\mathbb{R}^{m\times n}$ and $\xi$ represents a random variable, such as a training data point, sampled from an unknown data distribution $\mathcal{D}$.
+2 Revisiting the Muon Algorithms
+We consider the following optimization problem:
+$\min_{\mathbf{X}\in\mathbb{R}^{m\times n}}f(\mathbf{X}),\mathrm{~where~}f(\mathbf{X})=\mathbb{E}_{\xi\sim\mathcal{D}}[f(\mathbf{X};\xi)],$ (1)
+where $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ is the loss function, $\mathbf{X}$ denotes the decision variable, and $\xi$ represents a random variable(e.g., a training data sample) drawn from an unknown distribution $\mathcal{D}$.
 We assume that $f$ is differentiable and possibly nonconvex.
-3 Revisiting the Muon Algorithms
+The Muon optimizer begins by computing a momentum-based variance-reduced gradient update, similar in spirit to ADAM Kingma2014AdamAM, STORM Cutkosky2019MomentumBasedVR, and SGD with Nesterov momentum sutskever2013importance.
+The momentum term $\mathbf{M}_{t}$ is then projected onto the set of orthogonal matrices.
+This orthogonalization step equalizes the singular values, ensuring that no principal component direction dominates the optimization.
+Finally, the resulting scaled orthogonal matrix is used to update the model parameters.
+The Muon algorithm is summarized in Algorithm 1.
 Algorithm 1 Muon
-1:Require Initial parameters $\mathbf{X}_{0}\in\mathbb{R}^{m\times n}$, learning rate $\eta_{t}>0$, momentum parameter $\beta_{t}\in[0,1)$, variance-reduction parameter $\gamma\in[0,1]$, initial momentum $\mathbf{M}_{0}=0$, $\nabla f(\mathbf{X}_{0};\xi)=0$.
+1:Input: Initial parameters $\mathbf{X}_{0}\in\mathbb{R}^{m\times n}$, learning rate $\eta_{t}>0$, momentum parameter $\beta_{t}\in[0,1)$, variance-reduction parameter $\gamma\in[0,1]$, initial momentum $\mathbf{M}_{0}=0$, $\nabla f(\mathbf{X}_{0};\xi)=0$.
 2:for $t=1$ to $T$ do
@@
 10:end for
-11:Return Final parameters $\mathbf{X}_{t}$
-In practice, Algorithm 1 leverages a quintic Newton–Schulz iteration to efficiently compute a rank-preserving normalization of any input matrix $\mathbf{M}_{t}\in\mathbb{R}^{m\times n}$.
-This operation can be precisely defined via the Compact Singular Value Decomposition (SVD) of $\mathbf{M}_{t}$, expressed as $\mathbf{M}_{t}=\mathbf{U}_{r}\mathbf{\Sigma}_{r}\mathbf{V}_{r}^{\top}$, where $r$ is the rank of $\mathbf{M}_{t}$ and $\mathbf{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ contains only the $r$ positive singular values.
-The operator computes the matrix $\mathbf{O}_{t}\in\mathbb{R}^{m\times n}$ given by:
-$\mathbf{O}_{t}=\mathbf{U}_{r}\mathbf{V}_{r}^{\top}$
-This is equivalent to setting all positive singular values of $\mathbf{M}_{t}$ to 1 while leaving its null space unchanged, thus preserving the original rank $r$.
-It is worth noting that this formulation is algebraically equivalent to using the Moore-Penrose pseudoinverse: $((\mathbf{M}_{t}\mathbf{M}_{t}^{\top})^{+})^{1/2}\mathbf{M}_{t}$.
-Only in the specific case where $\mathbf{M}_{t}$ is full-rank does this result coincide with the solution to the Orthogonal Procrustes problem, $\min_{\mathbf{O}}\|\mathbf{O}-\mathbf{M}_{t}\|_{F}^{2}$ s.t.
-$\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}$.
-In the subsequent theoretical analysis, we assume that this operator can be solved exactly.
-Building upon this core operator, we investigate the following update rules:
+11:Output: Final parameters $\mathbf{X}_{t}$
+This orthogonalization step can be obtained from the Singular Value Decomposition (SVD) of $\mathbf{M}_{t}$:
+$\mathbf{M}_{t}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top},\qquad\mathbf{O}_{t}=\mathbf{U}\mathbf{V}^{\top}.$
+Equivalently, we have $\mathbf{O}_{t}=(\mathbf{M}_{t}\mathbf{M}_{t}^{\top})^{-1/2}\mathbf{M}_{t}$, which shows that orthogonalization reduces to computing an inverse square root rather than a full SVD.
+Since forming and decomposing $\mathbf{M}_{t}\mathbf{M}_{t}^{\top}$ remains costly, the popular Muon implementation jordan2024muon uses the quintic Newton–Schulz iteration to approximate the inverse square root.
+This recurrence converges in only a few steps (typically five), producing a numerically stable, rank-preserving orthogonalization of $\mathbf{M}_{t}$ that is nearly as accurate as SVD but far more efficient.
+Algorithm 1 incorporates two distinct strategies for momentum-based variance reduction, termed Muon-MVR1 and Muon-MVR2.
+These options present a fundamental trade-off between computational efficiency and theoretical rigor.
+MVR2 implements a principled variance reduction scheme at the cost of two gradient evaluations per step, while MVR1 serves as a computationally cheaper, single-gradient approximation.
+We detail both below.
 […]
-$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1}))$ (2)
-Further analysis reveals that when the parameters are set as $\beta_{t}=\mu$ and $\gamma=1-\beta_{t}$, the iterative process of Algorithm 1, subject to a rescaling of $\mathbf{M}_{t}$, becomes mathematically equivalent to the update scheme (3) of the Muon algorithm Jordan et al.
-(2024); Liu et al.
-(2025a).
-The Muon update is given by: $\mathbf{M}_{t}=\mu\mathbf{M}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t})+\mu(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1}))$
-$\mathbf{C}_{t}$ $=\mu\mathbf{C}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t})$ (3)
-$\mathbf{M}_{t}$ $=\mu\mathbf{C}_{t}+\nabla f(\mathbf{X}_{t};\xi_{t})$
-$\mathbf{O}_{t}$ $=\text{Newton-Schulz}(\mathbf{M}_{t})$
-$\mathbf{X}_{t+1}$ $=\mathbf{X}_{t}-\eta\mathbf{O}_{t}$
-Muon-MVR1 reduces to the EMA stochastic gradient estimator Kingma & Ba (2014) update Rule (4) as a special case when $\gamma=0$, and it can also be regarded as an approximation of MVR2.
-$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})$ (4)
+$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1})).$ (2)
+The primary advantage of this formulation is its computational efficiency, as it requires only one stochastic gradient evaluation per iteration. This update rule is flexible:
+(i) When $\gamma=0$, Rule (2) degenerates to the standard exponential moving average (EMA) of gradients (Rule (3)), a stochastic gradient estimator widely used in optimizers like Adam Kingma2014AdamAM.
+$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t}).$ (3)
+(ii) With the specific parameterization ($\beta_{t}=\mu,\gamma=1-\beta_{t}$), it can be shown that Rule (2) becomes mathematically equivalent to the update scheme of the Muon optimizer jordan2024muon; Liu2025MuonIS, given by:
+$\mathbf{C}_{t}$ $=\mu\mathbf{C}_{t-1}+\nabla f(\mathbf{X}_{t};\xi_{t}),$ (4)
+$\mathbf{M}_{t}$ $=\mu\mathbf{C}_{t}+\nabla f(\mathbf{X}_{t};\xi_{t}).$
 […]
-$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t}))$ (5)
-The key distinction between this rule and Rule (2) lies in the term subtracted from the current stochastic gradient, $\nabla f(\mathbf{X}_{t};\xi_{t})$.
-Specifically, Rule (5) subtracts the previous gradient computed using the current data sample $\xi_{t}$, namely $\nabla f(\mathbf{X}_{t-1};\xi_{t})$, a crucial modification designed to cancel the variance introduced by $\xi_{t}$ Yuan et al.
-(2024).
-In contrast, Rule (2) reuses the gradient based on the previous data sample $\xi_{t-1}$ (i.e., $\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$), effectively acting as a computationally efficient approximation of Rule (5).
-This approximation is justified because the difference between the two approaches is a noise term, $\Delta^{\text{Noise}}_{t-1}=\nabla f(\mathbf{X}_{t-1};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$.
-Under the standard assumption that stochastic gradients have a bounded variance, $\mathbb{E}_{\xi}[\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}]\leq\sigma^{2}$, the variance of this noise term is well-controlled and satisfies $\mathbb{E}[\|\Delta^{\text{Noise}}_{t-1}\|_{F}^{2}]\leq 2\sigma^{2}$.
-While Rule (2) offers the practical benefit of requiring only one gradient evaluation per step versus two for Rule (5), this structural difference leads to fundamentally different convergence properties.
-Our theoretical analysis demonstrates that the Muon-variant algorithm employing the stricter variance-reduction mechanism of Rule (5) achieves a convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ in Theorem 4.2.
+$\mathbf{M}_{t}=\beta_{t}\mathbf{M}_{t-1}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})+\gamma\cdot\beta_{t}\cdot(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t})).$ (5)
+The key distinction from MVR1 is the correction term.
+MVR2 subtracts the gradient computed on the previous parameters but with the current mini-batch, i.e., $\nabla f(\mathbf{X}_{t-1};\xi_{t})$.
+This modification is crucial as it is designed to directly cancel the variance introduced by the mini-batch $\xi_{t}$ Cutkosky2019MomentumBasedVR; Huang2021SUPERADAMFA; Yuan2024MARSUT.
+However, this theoretical benefit comes at the cost of requiring two gradient evaluations per step.
+MVR1 can be formally understood as a practical approximation of MVR2.
+The difference between their update rules is a single noise term, $\Delta^{\text{Noise}}_{t-1}=\nabla f(\mathbf{X}_{t-1};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t-1})$.
+Under the standard assumption of bounded variance ($\mathbb{E}_{\xi}[\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}]\leq\sigma^{2}$), the variance of this noise is well-controlled, satisfying $\mathbb{E}[\|\Delta^{\text{Noise}}_{t-1}\|_{F}^{2}]\leq 2\sigma^{2}$.
+While MVR1 is often sufficient in practice, this structural difference leads to fundamentally different theoretical guarantees.
+The principled variance cancellation in MVR2 allows our algorithm to achieve a superior iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$, as we will formally establish in Theorem 3.2.
 3 Convergence Analysis
 We begin our analysis by situating it within the established context of first-order stochastic optimization.
-The $\mathcal{O}(T^{-1/4})$ convergence rate is a well-known bottleneck for methods like SGD, and recent work by Shen et al.
-(2025) confirmed this limitation for the standard Muon algorithm.
-Accordingly, our first step in this section is to analyze the Muon-MVR1 variant and formally reaffirm this baseline convergence rate in Theorem 4.1.
+The $\mathcal{O}(T^{-1/4})$ iteration complexity is a well-known bottleneck for methods like SGD, and recent work by shen2025convergence confirmed this limitation for the standard Muon algorithm.
+Accordingly, our first step in this section is to analyze the Muon-MVR1 variant and formally reaffirm this baseline iteration complexity in Theorem 3.1.
 While the rate itself is standard, our primary contribution here lies in the novel and unified analytical framework we develop to prove it, as this framework will be instrumental for subsequent results.
@@
 Specifically, there exists a constant $\sigma>0$ such that for all $\mathbf{X}\in\mathbb{R}^{m\times n}$, the following holds: $\mathbb{E}[\nabla f(\mathbf{X};\xi)]=\nabla f(\mathbf{X})$ and $\mathbb{E}\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}\leq\sigma^{2}$.
-These assumptions are quite common Zhou et al. (2018); Chen et al. (2018); Huang et al. (2021); Guo et al. (2021); Li et al. (2023); Wang et al. (2023); Xie et al. (2024); Yuan et al. (2024).
+These assumptions are quite common Zhou2018OnTC; Chen2018OnTC; Huang2021SUPERADAMFA; Guo2021ANC; Li2023ConvergenceOA; Wang2023ClosingTG; 10586270; Yuan2024MARSUT.
 3.1 Erogdic Convergence of Muon
+In this subsection, we establish the ergodic convergence of Muon-MVR1 and Muon-MVR2.
 3.1.1 Option MVR1
+We begin our analysis with Option MVR1, a straightforward implementation of one-batch momentum-based variance reduction.
+The following theorem establishes its ergodic convergence rate, demonstrating that the algorithm converges to a stationary point at a rate of $\tilde{\mathcal{O}}(T^{-1/4})$ for specific choices of learning rate and momentum schedules.
 Theorem 3.1.
-Suppose Assumptions 4.1, 4.2, and 4.4 hold. Consider Algorithm 1 with a learning rate of $\eta_{t}=t^{-3/4}$. The expected average squared norm of the gradient is bounded for the following options:
+Suppose Assumptions 3.1, 3.2, and 3.4 hold. Consider Algorithm 1 with a learning rate of $\eta_{t}=t^{-3/4}$. The expected average squared norm of the gradient is bounded for the following options:
   1. 1.
@@
 The size of this neighborhood is determined by the stochastic gradient variance, the learning rate, and the momentum parameter.
-Notably, this convergence bound is free of any non-vanishing additive error term that depends on the dimension $n$ Sato et al.
-(2025).
+Notably, this convergence bound is free of any non-vanishing additive error term that depends on the dimension $n$ Sato2025ConvergenceBA.
 We absorb the dimensional dependence into the $\mathcal{O}$-notation to define $\tilde{\mathcal{O}}$, which yields the following complexity for the algorithm:
-$\min_{t=1,\dots,T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\tilde{\mathcal{O}}(T^{-1/4})$
+$\min_{t=1,\dots,T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\tilde{\mathcal{O}}(T^{-1/4}).$
 3.1.2 Option MVR2
-While prior work has established a convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ for variance-reduction in non-convex settings for various algorithms, such as SGD Fang et al.
-(2018); Cutkosky & Orabona (2019); Zhou et al.
-(2020) and Adam Huang et al.
-(2021); Yuan et al.
-(2024), the theoretical convergence properties of this technique when applied to the Muon optimizer have remained an open question.
-We bridge this theoretical gap in Theorem 4.2, which rigorously proves that Option MVR2 of Algorithm 1 (Muon-MVR2) achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ convergence rate in the general non-convex setting.
+While prior work has established a iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$ for variance-reduction in non-convex settings for various algorithms, such as SGD fang2018spider; Cutkosky2019MomentumBasedVR; zhou2020stochastic and Adam Huang2021SUPERADAMFA; Yuan2024MARSUT, the theoretical convergence properties of this technique when applied to the Muon optimizer have remained an open question.
+We bridge this theoretical gap in Theorem 3.2, which rigorously proves that Option MVR2 of Algorithm 1 (Muon-MVR2) achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ iteration complexity in the general non-convex setting.
 Theorem 3.2.
-Under Assumptions 4.1, 4.3, and 4.4, consider Algorithm 1 with Option MVR2 and update rule (5) parameters set as $\beta_{t}=1-\eta_{t}$, $\eta_{t}=t^{-2/3}$ for $t\geq 1$, and $\gamma=1$.
+Under Assumptions 3.1, 3.3, and 3.4, consider Algorithm 1 with Option MVR2 and update rule (5) parameters set as $\beta_{t}=1-\eta_{t}$, $\eta_{t}=t^{-2/3}$ for $t\geq 1$, and $\gamma=1$.
 Then, the following bounds hold:
-[…]
+$\begin{aligned} \frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}&\leq\frac{\sqrt{4\sigma^{2}+(16L^{2}n+8\sigma^{2})(1+\ln T)}}{T^{1/3}}+\frac{G}{T^{1/3}},\end{aligned}$
 where
@@
 Theorem 3.2 establishes a significant theoretical advantage.
-In contrast to prior work Sfyraki & Wang (2025) that requires a computationally expensive growing batch size ($b=\mathcal{O}(T^{1/3})$) to control gradient variance, our result shows that Muon-MVR2 achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ convergence rate in a more practical, fixed mini-batch setting.
+In contrast to prior work sfyraki2025lions that requires a computationally expensive growing batch size ($b=\mathcal{O}(T^{1/3})$) to control gradient variance, our result shows that Muon-MVR2 achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ iteration complexity in a more practical, fixed mini-batch setting.
 This is made possible by the specific hyperparameter schedule, where the learning rate $\eta_{t}=t^{-2/3}$ and the momentum parameter $\beta_{t}=1-\eta_{t}$ are critically coupled.
@@
 Remark 3.3.
-As noted in Yuan et al.
-(2024), a more sophisticated, adaptive setting for the variance-reduction parameter $\gamma$ can be employed.
+As noted in Yuan2024MARSUT, a more sophisticated, adaptive setting for the variance-reduction parameter $\gamma$ can be employed.
 Specifically, by setting $\gamma=\gamma_{t}=1-\frac{A_{t}}{\beta_{t}}$, where $A_{t}$ is defined in Lemma C.2, a key term in the analysis becomes $P_{t}=-\mathbb{E}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t})\|_{F}^{2}\cdot A_{t}^{2}$, which is strictly negative.
@@
 Remark 3.4.
-The PL condition has been widely employed in the convergence analysis of various first-order algorithms Karimi et al.
-(2016); Xie et al.
-(2020a); Li & Li (2022), though typically under restricted settings.
+The PL condition has been widely employed in the convergence analysis of various first-order algorithms karimi2016linear; xie2020linear; li2022simple, though typically under restricted settings.
 Note that if $f(\cdot)$ is strongly convex, then it is necessarily convex and satisfies the PL condition.
@@
 Theorem 3.3.
-Suppose Assumptions 4.1, 4.2, 4.4, and 4.5 hold.
+Suppose Assumptions 3.1, 3.2, 3.4, and 3.5 hold.
 Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with a step size of $\eta_{t}=t^{-3/4}$.
-For either of the following options:
+We analyze the following two MVR1 schemes:
   1. 1.
-    For Option MVR1 ($\gamma=0$), using the update rule (4) with momentum $\beta_{t}=1-t^{-1/2}$,
+    Scheme 1 ($\gamma=0$): Using the update rule (3) with momentum $\beta_{t}=1-t^{-1/2}$. For this case, we define the constant $\mathcal{A}_{1}$ as:
+    $\mathcal{A}_{1}=6L^{-1}\sigma^{2}+(8\sqrt{2}+2)Ln+L.$
   2. 2.
-    For Option MVR1 ($\gamma_{t}=t^{-1/2}$), using the update rule (2) with parameters $\beta_{t}=1-(t+1)^{-1/2}$,
-the expected suboptimality gap after $T$ iterations is bounded by:
-$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right).$
+    Scheme 2 ($\gamma_{t}=t^{-1/2}$): Using the update rule (2) with parameters $\beta_{t}=1-(t+1)^{-1/2}$. For this case, we define the constant $\mathcal{A}_{2}$ as:
+    $\mathcal{A}_{2}=14L^{-1}\sigma^{2}+(16\sqrt{2}+2)Ln+L.$
+Then, for either scheme $i\in\{1,2\}$, there exists a constant $T_{0}=2e^{4}$ such that for all iterations $T\geq T_{0}$, the expected suboptimality gap is bounded as follows:
+$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{\mathcal{A}_{i}^{2}}{8\mu}\cdot\frac{(\ln T)^{2}}{T^{1/2}}.$
 Theorem 3.4.
-Suppose Assumptions 4.1, 4.2, 4.4, and 4.5 hold.
+Suppose Assumptions 3.1, 3.2, 3.4, and 3.5 hold.
 Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with Option MVR2 (see update rule (5)).
-By setting the parameters to $\eta_{t}=t^{-2/3}$, $\beta_{t}=1-t^{-2/3}$, and $\gamma=1$, the expected suboptimality gap after $T$ iterations is bounded by:
-$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right).$
+By setting the parameters to $\eta_{t}=t^{-2/3}$, $\beta_{t}=1-t^{-2/3}$, and $\gamma=1$, we define the constant $\mathcal{A}_{3}$ as :
+$\mathcal{A}_{3}=20L^{-1}\sigma^{2}+34Ln+L.$
+Then, there exists a constant $T_{0}=2e^{3}$ such that for all iterations $T\geq T_{0}$, the expected suboptimality gap is bounded as follows:
+$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{2\mathcal{A}_{3}^{2}}{9\mu}\cdot\frac{(\ln T)^{2}}{T^{2/3}}.$
+Detailed proofs for Theorems 3.3 and 3.4 are provided in Appendix F and G, respectively.
 Remark 3.5.
-The convergence bounds in Theorems 4.3, and 4.4 are asymptotic, holding for any number of iterations $T$ exceeding a sufficiently large threshold $T_{A}$.
-The existence of this threshold is established in our proofs and depends on the selection of a helper constant $A$.
-The specific conditions on $A$ and definitions of $\mathcal{A}$ are as follows:
-  • For Theorem 4.3 (MVR1 with $\gamma=0$): $A>\frac{\mathcal{A}}{8C}$ with $\mathcal{A}=6L^{-1}\sigma^{2}+8\sqrt{2}Ln+2Ln+L$.
-  • For Theorem 4.3 (MVR1): $A>\frac{\mathcal{A}}{8C}$ with $\mathcal{A}=14L^{-1}\sigma^{2}+16\sqrt{2}Ln+2Ln+L$.
-  • For Theorem 4.4 (MVR2): $A>\frac{\mathcal{A}}{6C}$ with $\mathcal{A}=20L^{-1}\sigma^{2}+34Ln+L$.
-In all cases, the constant $C$ is defined as $C=\frac{\sqrt{2\mu}}{2}$. The proofs of Theorems 4.3, and 4.4 are provided in Appendix F, and G, respectively.
-Remark 4.6.
-Theorem 4.4 achieves a faster convergence rate of $\tilde{\mathcal{O}}(T^{-2/3})$ than the $\tilde{\mathcal{O}}(T^{-1/2})$ in Theorem 4.3, owing to its more effective variance reduction.
-While our rates are not as fast as the $\mathcal{O}(1/T)$ rate for Adam He et al.
-(2023), our analysis is more general, as it does not require the bounded stochastic gradient assumption used therein.
-Closing this rate gap under our relaxed assumption is an important direction for future work.
-Remark 4.7.
-Our non-ergodic convergence proofs are unified by a key recursive inequality for a Lyapunov function $\Delta_{t}=\mathbb{E}[f(\mathbf{X}_{t})]-f^{*}$:
-$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}.$
-The final convergence rate is dictated by the exponent $p\in(0,1)$.
-We establish this with $p=3/4$ for Theorem 4.3.
-For Theorem 4.4, superior variance control allows proving the inequality with a tighter $p=2/3$, leading to the faster rate.
+Our non-ergodic convergence proofs are unified by a recursive inequality on the potential function $\Delta_{t}=\mathbb{E}[f(\mathbf{X}_{t})]-f^{*}$:
+$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t},$
+where $\sum_{i=1}^{t}\Gamma_{t}\leq\mathcal{O}(\ln t)$.
+The exponent $p\in(0,1)$ dictates the final convergence rate.
+For Theorem 3.3, we establish this inequality with $p=3/4$, which leads to a rate of $\tilde{\mathcal{O}}(T^{-1/2})$.
+The superior variance control in Theorem 3.4 enables a tighter analysis with $p=2/3$, which in turn yields the accelerated convergence rate of $\tilde{\mathcal{O}}(T^{-2/3})$.
 4 Experiments
@@
 However, its computational overhead can be prohibitive in large-scale settings, as shown in Figure 1c.
-We note that prior analysis in Sfyraki & Wang (2025); Sato et al.
-(2025) suggests that increasing batch size can mitigate the impact of stochastic noise on convergence properties.
+We note that prior analysis in sfyraki2025lions; Sato2025ConvergenceBA suggests that increasing batch size can mitigate the impact of stochastic noise on convergence properties.
 Given that mini-batch training is the standard paradigm in practice, a trade-off between theoretical exactness and computational feasibility is often inherent.
-Therefore, echoing the sentiment in Yuan et al.
-(2024), we posit that the approximate version serves as a pragmatic alternative when computational efficiency is a primary concern, as it incurs only a minimal degradation in performance.
+Therefore, echoing the sentiment in Yuan2024MARSUT, we posit that the approximate version serves as a pragmatic alternative when computational efficiency is a primary concern, as it incurs only a minimal degradation in performance.
 Conversely, in scenarios where maximizing validation accuracy is paramount, the precise formulation remains the recommended approach.
@@
 5.1 The Evolution of Optimization Algorithms
-One of the cornerstones of deep learning is first-order optimization methods.
-In early research, the Momentum method proposed by Nesterov (1983) accelerates the convergence process by accumulating historical gradient information.
-Subsequently, the RMSprop algorithm introduced by Hinton et al.
-(2012) pioneered the concept of adaptive learning rates, allowing different parameters to have independent update step sizes.
-The Adam optimizer, proposed by Kingma & Ba (2014), integrated these two ideas and has become one of the most widely used optimizers in the deep learning field by adaptively adjusting the first and second moments of the gradients.
-Building upon Adam, the academic community has proposed a series of improvements to address its limitations in various scenarios.
-To enhance the stability of the optimization process, Reddi et al.
-(2018) proposed AMSGrad, which uses the maximum of historical second moments to constrain the learning rate update.
-Dozat (2016) integrated Nesterov momentum into Adam to design NAdam, aiming for superior performance.
-To address issues in Adam concerning weight decay and learning rate adjustment, Loshchilov (2017) developed AdamW, which improves the effectiveness of L2 regularization by decoupling weight decay.
-Luo et al.
-(2019) designed AdaBound, which sets dynamic bounds on the learning rate to avoid excessively large or small update steps.
-Furthermore, to mitigate variance fluctuations in the early stages of training, Liu et al.
-(2019) proposed RAdam to rectify the variance estimate.
-Zhuang et al.
-(2020) proposed AdaBelief, which adjusts the second moment by calculating the deviation between the gradient and its moving average, aiming to enhance the model’s generalization ability.
-In recent years, more novel Adam variants have been proposed.
-For instance, Xie et al.
-(2020b), from a dynamical systems perspective, constructed the Adai framework to accelerate training and improve the quality of convergence points by decoupling the effects of the adaptive learning rate and momentum.
-Xie et al.
-(2024) proposed Adan, which introduces a novel Nesterov momentum estimation method to accelerate convergence without incurring additional gradient computation overhead.
-The latest works, such as Liang et al.
-(2024), introduce a Cautious masking strategy to enhance optimization efficiency.
-Yuan et al.
-(2024) proposed MARS to further improve the variance-reduction strategy in Adam.
-Beyond direct improvements to Adam, researchers have also explored other efficient optimization paradigms.
-A significant branch among these is Shampoo-type methods based on preconditioning with parameter curvature.
-Shampoo, proposed by Gupta et al.
-(2018), is one of the pioneering works in this area.
-Inspired by its framework, Muon Jordan et al.
-(2024); Liu et al.
-(2025a) was proposed.
-This method adapts to parameter curvature by orthogonalizing gradient momentum via the Newton-Schulz iteration, thereby improving convergence dynamics.
-Recently, several adaptive variants of Muon have also emerged.
-AdaMuon Si et al.
-(2025) incorporates element-wise adaptivity but lacks theoretical convergence guarantees.
-For efficient training of large language models, COSMOS Liu et al.
-(2025b) integrates ideas from SOAP Vyas et al.
-(2025) and Muon, demonstrating stability and memory advantages in practice but similarly lacks a convergence proof.
-In the literature, other related preconditioning methods include ASGO An et al.
-(2025), PolarGrad Lau et al.
-(2025), and AdaGO Zhang et al.
-(2025), which introduces the Adagrad-Norm step size Ward et al.
-(2020) into a simplified version of Muon.
-Meanwhile, other high-performing optimizers not belonging to the Shampoo family also warrant attention, such as Sophia Liu et al.
-(2023), which improves second-moment estimation through efficient diagonal Hessian approximation and coordinate clipping, and Lion Chen et al.
-(2023), a lightweight optimizer that only tracks momentum and uses the sign function to normalize updates.
+The advancement of deep learning relies on first-order optimization methods.
+Nesterov1983AMF proposed the Momentum method, which leverages historical gradients to accelerate convergence.
+Subsequently, hinton2012lecture introduced RMSprop, enabling per-parameter adaptive learning rates.
+Kingma2014AdamAM then integrated these ideas in Adam, an optimizer that adapts using the first and second moments of the gradients, establishing it as a standard choice in the field.
+Numerous variants have been proposed to address the limitations of Adam.
+Reddi2018OnTC introduced AMSGrad to ensure a non-increasing learning rate, while  dozat2016incorporating created NAdam by incorporating Nesterov momentum.
+To improve regularization,  loshchilov2017decoupled developed AdamW, which decouples weight decay from the optimization step.
+Other methods focus on controlling the learning rate and variance:  Luo2019AdaptiveGM proposed AdaBound to dynamically clip learning rates;  Liu2019OnTV designed RAdam to rectify variance estimates in early training; and  Zhuang2020AdaBeliefOA created AdaBelief to adapt step sizes based on gradient belief.
+Recent developments continue this trend, such as  Xie2020AdaptiveID which decouples adaptation from momentum in the Adai framework, and  10586270 which introduced Adan with a novel Nesterov momentum estimation.
+The latest works from  Liang2024CautiousOI and  Yuan2024MARSUT further enhance efficiency with new masking and variance-reduction strategies, respectively.
+Additionally, the MGUP strategy proposed by chang2025mgup smooths zero masks into small non-zero values to alleviate potential non-convergence issues that the Cautious mask may cause in Adam.
+Beyond Adam variants, research has explored other paradigms such as preconditioning methods that use parameter curvature.
+Gupta2018ShampooPS pioneered this direction with the Shampoo optimizer.
+Building on this work,  jordan2024muon proposed Muon, which adapts to curvature by orthogonalizing gradient momentum.
+Subsequent variants emerged, such as AdaMuon si2025adamuon which adds element-wise adaptivity, and COSMOS liu2025cosmos which integrates ideas from SOAP vyas2025soap for large model training.
+While these methods showed practical benefits, they often lacked convergence proofs.
+To bridge this theoretical gap for LMO-based methods, Gluon riabinin2025gluon introduces a novel layer-wise smoothness assumption, providing convergence guarantees that align with the practical implementations of optimizers like Muon and Scion pethick2025training.
+In the literature, other related preconditioning methods include ASGO An2025ASGOAS, PolarGrad lau2025polargrad, and AdaGO Zhang2025AdaGradMM, which introduces the Adagrad-Norm step size ward2020adagrad into a simplified version of Muon.
+Meanwhile, other high-performing optimizers not belonging to the Shampoo family also warrant attention, such as Sophia Liu2023SophiaAS, which improves second-moment estimation through efficient diagonal Hessian approximation and coordinate clipping, and Lion Chen2023SymbolicDO, a lightweight optimizer that only tracks momentum and uses the sign function to normalize updates.
 5.2 Analysis of Muon
-Recent studies have elucidated the underlying mechanisms and convergence properties of the Muon optimizer from multiple theoretical perspectives.
-A central viewpoint reinterprets modern optimizers as steepest descent methods under norm constraints.
-Although this analysis does not directly incorporate momentum, it offers valuable insights into the update directions of optimizers like Muon and Adam Bernstein & Newhouse (2024).
-Building on this, subsequent research has established a theoretical link between the Muon algorithm with weight decay and the stochastic Frank-Wolfe method, specifically identifying Muon as an instance of stochastic Frank-Wolfe under a spectral norm constraint Sfyraki & Wang (2025).
-This provides a foundational understanding of its behavior from a constrained optimization perspective.
-Other work has characterized Muon as a special case of Lion-$\mathcal{K}$.
-Separately, its convergence to a Karush-Kuhn-Tucker (KKT) point has been proven in both deterministic and stochastic gradient settings Chen et al.
-(2025).
-These theories, together with research on stochastic conditional gradient methods from a norm-constrained viewpoint, form the theoretical underpinnings for Muon Pethick et al.
-(2025).
-Beyond its theoretical framework, Muon’s implicit bias has garnered significant attention.
-It has been shown that Muon tends to converge to a max-margin solution with respect to the spectral norm of the classifier’s weight matrix, highlighting a distinct difference in implicit regularization and generalization preference compared to Adam Fan et al.
-(2025).
-A unifying preconditioning perspective explains this phenomenon through matrix polar decomposition, which explicitly distinguishes between curvature anisotropy and gradient anisotropy in optimization problems, thereby clarifying the behavioral differences between Muon and Adam when handling different types of parameters Lau et al.
-(2025).
-Regarding convergence analysis, while rigorous proofs and convergence rates for Muon in non-convex settings have been provided under various smoothness assumptions (e.g., Frobenius-norm or spectral-norm Lipschitz smoothness) Shen et al.
-(2025), the validity of some early analyses has been questioned.
-For instance, certain convergence results are considered dubious due to the use of incorrect mathematical inequalities in their proofs Li & Hong (2025).
-Furthermore, other analyses, despite proving the convergence of several Muon variants, have indicated slow convergence rates or the need for stringent conditions to reach a desired stationary point Sato et al.
-(2025).
-On a practical level, particularly in the pre-training of large-scale language models (LLMs), weight decay has been shown to be an indispensable component.
-Concurrently, methods for effectively transferring learning rates from Adam to Muon, based on RMS analysis, have been proposed to enhance its scalability and utility Liu et al.
-(2025a).
+Recent theoretical analyses have clarified the mechanisms and convergence of the Muon optimizer.
+One central perspective frames modern optimizers as steepest descent methods under specific norm constraints; while this view does not directly account for momentum, it offers crucial insights into the update directions of optimizers like Muon and Adam bernstein2024old.
+Building on this, subsequent work has linked Muon with weight decay to the stochastic Frank-Wolfe method, identifying it as an instance operating under a spectral norm constraint sfyraki2025lions.
+From a constrained optimization viewpoint, Muon has also been characterized as a special case of Lion-$\mathcal{K}$, with its convergence to a KKT point proven in both deterministic and stochastic settings chen2025muon.
+These analyses, alongside broader research on norm-constrained stochastic conditional gradient methods, form Muon’s theoretical underpinnings pethick2025training.
+Beyond its formal framework, Muon’s implicit bias has become a key area of investigation.
+It has been demonstrated that Muon tends to converge toward max-margin solutions with respect to the spectral norm of the weight matrix, revealing a distinct implicit regularization and generalization preference compared to Adam fan2025implicit.
+A unifying preconditioning perspective, based on matrix polar decomposition, explains this behavior by distinguishing between curvature and gradient anisotropy.
+This decomposition clarifies the differing ways Muon and Adam handle various parameter types lau2025polargrad.
+The convergence analysis of Muon has also seen active development.
+Although rigorous proofs and iteration complexity in non-convex settings have been established under various smoothness assumptions shen2025convergence, the validity of some early results has been challenged.
+For instance, certain proofs were found to rely on incorrect mathematical inequalities, casting doubt on their conclusions Li2025ANO.
+Other analyses, while confirming the convergence of several Muon variants, report slow convergence rates or the need for stringent conditions to reach a stationary point Sato2025ConvergenceBA.
+On a practical level, especially for pre-training large-scale language models (LLMs), weight decay has been identified as an indispensable component for Muon.
+To enhance its scalability and utility, methods based on RMS analysis have also been proposed to effectively transfer learning rates from Adam to Muon Liu2025MuonIS.
 6 Conclusion
-In this work, we established a rigorous theoretical foundation for the Muon optimizer, addressing the significant gap between its empirical success and its formal analysis.
-We introduced Muon-MVR2, a variance-reduced variant, and provided the first proof that it achieves an optimal convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ in stochastic non-convex settings, matching the theoretical lower bound Arjevani et al.
-(2023).
-Furthermore, we demonstrated that under the Polyak-Łojasiewicz condition, the Muon-MVR1, and Muon-MVR2 variants all attain sublinear non-ergodic convergence.
-Our theoretical results are validated by extensive experiments on CIFAR-10 and C4 benchmarks, which confirm the practical acceleration and superior performance of Muon-MVR2 over its standard counterpart and other widely-used optimizers.
-This research solidifies the Muon framework by providing robust theoretical guarantees and a practically effective new variant for training deep learning models.
+In this work, we establish a rigorous theoretical foundation for the Muon optimizer, addressing the gap between its empirical success and formal analysis.
+We analyze two momentum-based variance-reduced variants of Muon: a one-batch version (Muon-MVR1) and a two-batch version (Muon-MVR2).
+We provide the first proof that Muon-MVR2 achieves the optimal iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$ in stochastic nonconvex settings, matching the theoretical lower bound arjevani2023lower.
+Furthermore, we show that under the Polyak–Łojasiewicz condition, Muon-MVR1 and Muon-MVR2 attain sublinear nonergodic convergence rates of $\tilde{\mathcal{O}}(T^{-1/2})$ and $\tilde{\mathcal{O}}(T^{-2/3})$, respectively.
+Our theoretical results are validated by extensive experiments on the CIFAR-10 and C4 benchmarks, which confirm the practical acceleration and superior performance of Muon-MVR2 over its standard counterpart and other widely used optimizers.
+Overall, this research strengthens the Muon framework by providing robust theoretical guarantees and a practically effective new variant for deep learning training.
 Appendix
@@
 For Algorithm 1, choosing an arbitrary parameter $\alpha>0$, we have the following inequality:
-$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta^{2}n}{2}$
+$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta^{2}n}{2}.$
 Proof.
@@
 $\leq f(\mathbf{X}_{t})-\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
-$\leq f(\mathbf{X}_{t})-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle-\eta_{t}\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
+$\leq f(\mathbf{X}_{t})-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle-\eta_{t}\langle\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}.$
 Now we bound the three terms on the right-hand side respectively:
 Descent Term: According to the definition of $\mathbf{O}_{t}$ and the norm property $\|\mathbf{M}_{t}\|_{*}\geq\|\mathbf{M}_{t}\|_{F}$, we have:
-$\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}$
+$\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}.$
 Cross Term: This is the key to eliminating the dimension-dependent error.
@@
 $\leq\eta_{t}\left(\frac{\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{1}{2\alpha}\|\mathbf{O}_{t}\|_{F}^{2}\right)$
-$\leq\eta_{t}\left(\frac{\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{n}{2\alpha}\right)$
+$\leq\eta_{t}\left(\frac{\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{n}{2\alpha}\right).$
 Quadratic Term:
-$\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}\leq\frac{L\eta_{t}^{2}n}{2}$
+$\frac{L\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}\leq\frac{L\eta_{t}^{2}n}{2}.$
 Substituting these three bounds into the inequality for $f(\mathbf{X}_{t+1})$:
-$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta_{t}^{2}n}{2}$
+$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}\alpha}{2}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}n}{2\alpha}+\frac{L\eta_{t}^{2}n}{2}.$
 This completes the proof. ∎
@@
 For Algorithm 1 option MVR1 ($\gamma=0$), the accumulated error between the momentum term and the true gradient is bounded:
-$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}$
+$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
 Proof.
@@
 $\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
-$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
+$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}.$
 According to Assumption 3.4. Taking the expectation of its squared norm, and using the unbiasedness and independence of the stochastic gradient, we obtain:
@@
 $+\beta_{t+1}^{2}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}]$
-$+2\beta_{t+1}^{2}\mathbb{E}[\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle]$
+$+2\beta_{t+1}^{2}\mathbb{E}[\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle].$
 Applying Young’s inequality with a parameter ($ab\leq\frac{\epsilon}{2}a^{2}+\frac{1}{2\epsilon}b^{2}$), we have
@@
 $\quad+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$
-$\quad+(1-\beta_{t+1})^{2}\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$
+$\quad+(1-\beta_{t+1})^{2}\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right].$
 According to Assumption 3.2,
@@
 $=L^{2}\eta_{t}^{2}\|\mathbf{O}_{t}\|_{F}^{2}$
-$\leq L^{2}\eta_{t}^{2}n$
+$\leq L^{2}\eta_{t}^{2}n.$
 Therefore:
@@
 Then, by letting $\epsilon:=\frac{1-\beta_{t}}{\beta_{t}}$, we have
-$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}t\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}$ (6)
+$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$ (6)
 A.3 Lemma A.3
@@
 Proof.
+We derive the following inequalities:
 $\alpha_{t}E_{t}-c\left(E_{t}-E_{t+1}+A_{t+1}\right)$
@@
 $=E_{t}\cdot(t+1)^{-p}\cdot\left(\left(\frac{t}{t+1}\right)^{-p}-c\right)$
-$\stackrel{{\scriptstyle(\circ)}}{{\leq}}E_{t}\cdot(t+1)^{-p}\cdot(2-c),$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}0$
+$\stackrel{{\scriptstyle(\circ)}}{{\leq}}E_{t}\cdot(t+1)^{-p}\cdot(2-c)$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}0,$
 where $(\bullet)$ follows from $E_{t+1}\leq(1-\alpha_{t+1})E_{t}+A_{t+1}$; $(\circ)$ is due to $(\frac{t}{t+1})^{-p}\leq 2^{p}\leq 2$; $(\star)$ is due to our choice $c=2$. ∎
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}^{2/3}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\eta_{t}^{4/3}Ln}{2}+\frac{L\eta_{t}^{2}n}{2}$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}^{2/3}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+Ln\eta_{t}^{4/3}$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\eta_{t}^{2/3}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+Ln\eta_{t}^{4/3},$
 where $(\circ)$ by setting $\alpha=\frac{1}{\eta_{t}^{1/3}L}$; $(\star)$ follows from $\eta_{t}\leq 1$, we have $L\leq L/\eta_{t}^{1/3}$.
@@
 $\stackrel{{\scriptstyle(\bullet)}}{{=}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]$
-$+\underbrace{\frac{L\eta_{t}^{4/3}}{2}+\left(\frac{\eta_{t}^{2/3}}{2L}+\frac{\eta_{t}^{2/3}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}}_{\Gamma_{t}}$
+$+\underbrace{\frac{L\eta_{t}^{4/3}}{2}+\left(\frac{\eta_{t}^{2/3}}{2L}+\frac{\eta_{t}^{2/3}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{4/3}}_{\Gamma_{t}},$
 where $(\circ)$ follows from the reverse triangle inequality $\|\mathbf{M}_{t}\|_{F}\leq\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}-\|\nabla f(\mathbf{X}_{t})\|_{F}$; $(\star)$ applies Young’s inequality to the term $\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}]$; and $(\bullet)$ collects the residual terms into $\Gamma_{t}$ and sets $\epsilon=\eta_{t}^{2/3}/L$.
@@
 $\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$ $\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}$
-$\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{L^{2}\eta_{t}^{2}n}{1-\beta_{t+1}}+(1-\beta_{t+1})^{2}\sigma^{2}$
+$\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{L^{2}\eta_{t}^{2}n}{1-\beta_{t+1}}+(1-\beta_{t+1})^{2}\sigma^{2}.$
 Let $\mathbf{S}_{t+1}=\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})$. Thus, setting $\alpha_{t+1}=(t+1)^{-1/2}$, we observe the following relationship for $t\geq 1$:
-$\frac{\eta_{t}^{2}}{1-\beta_{t+1}}=\frac{t^{-3/2}}{(t+1)^{-1/2}}=\frac{\sqrt{t+1}}{t^{3/2}}\leq\frac{2\sqrt{2}}{t+1}$
+$\frac{\eta_{t}^{2}}{1-\beta_{t+1}}=\frac{t^{-3/2}}{(t+1)^{-1/2}}=\frac{\sqrt{t+1}}{t^{3/2}}\leq\frac{2\sqrt{2}}{t+1}.$
 This allows us to bound the expectation as follows:
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\alpha_{t+1}^{2}(2\sqrt{2}L^{2}n+\sigma^{2})$
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\alpha_{t+1}^{2}(2\sqrt{2}L^{2}n+\sigma^{2}).$
 According to Lemma A.3, by letting $A_{t+1}=\alpha_{t+1}^{2}(2\sqrt{2}L^{2}n+\sigma^{2})$, we have
-$\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1})$
+$\alpha_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}).$
 Furthermore, since
 $\mathbb{E}\|\mathbf{S}_{1}\|$ $=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\mathbf{M}_{1}\|_{F}^{2}=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-(1-\beta_{1})\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}$
-$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{1}=0)$
+$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{1}=0).$
 It follows that
@@
 $\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t+1}$
-$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
+$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1).$
 Thus,
 $\Gamma_{t}$ $=\frac{\eta_{t}^{2/3}}{L}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\eta_{t}^{4/3}$
-$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}$
+$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}.$
 Next, we define $A_{1}=4\sqrt{2}Ln+Ln+2L^{-1}\sigma^{2}+L/2$ and $A_{2}=4\sqrt{2}Ln+Ln+4L^{-1}\sigma^{2}+L/2$.
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)}{L}+(L/2+Ln)(\ln T+1)$
-$\leq A_{1}\ln T+A_{2}$
+$\leq A_{1}\ln T+A_{2},$
 where $(\circ)$ due to inequality (8).
@@
 $\leq\frac{f(\mathbf{X}_{1})-f^{*}}{T^{1/4}}+\frac{A_{1}\ln T+A_{2}}{T^{1/4}}$
-$=\tilde{\mathcal{O}}(T^{-1/4})$
+$=\tilde{\mathcal{O}}(T^{-1/4}).$
 Case 2: $\gamma\neq 0$.
@@
 $\mathbf{C}_{t}$ $=\beta_{t-1}\mathbf{C}_{t-1}+(1-\beta_{t-1})\nabla f(\mathbf{X}_{t};\xi_{t})$ (10)
-$\mathbf{M}_{t}$ $=\beta_{t}\mathbf{C}_{t}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})$
+$\mathbf{M}_{t}$ $=\beta_{t}\mathbf{C}_{t}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t}).$
 In this case, then we have
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}\beta_{t-1}\mathbb{E}\|\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1})\|_{F}^{2}+\frac{\beta_{t-1}^{2}L\eta_{t-1}^{2}n}{1-\beta_{t-1}}+(1-\beta_{t-1})^{2}\sigma^{2}$
-$=(1-\alpha_{t})\mathbb{E}\|\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1})\|_{F}^{2}+\alpha_{t}^{2}(2\sqrt{2}Ln+\sigma^{2})$
+$=(1-\alpha_{t})\mathbb{E}\|\mathbf{C}_{t-1}-\nabla f(\mathbf{X}_{t-1})\|_{F}^{2}+\alpha_{t}^{2}(2\sqrt{2}Ln+\sigma^{2}),$
 where ($\circ$) follows from Lemma A.2.
@@
 Using the conclusion of Lemma A.3, we have
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+A_{t+1}$
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\alpha_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+A_{t+1}.$
 Furthermore, since
 $\mathbb{E}\|\mathbf{S}_{1}\|$ $=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\mathbf{C}_{1}\|_{F}^{2}=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-(1-\beta_{0})\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}$
-$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{0}=0)$
+$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}\quad\quad\quad(\text{since }\beta_{0}=0).$
 Thus, as in inequality (8), we have
@@
 $\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t+1}$
-$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
+$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1).$
 Then,
 $\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}$ $=\|\beta_{t}\mathbf{C}_{t}+(1-\beta_{t})\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}$
-$=\|\beta_{t}(\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t}))+(1-\beta_{t})(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t}))\|_{F}$
+$=\|\beta_{t}(\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t}))+(1-\beta_{t})(\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t}))\|_{F}.$
 From this, we can bound the squared norm using the inequality $\|\mathbf{A}+\mathbf{B}\|_{F}^{2}\leq 2\|\mathbf{A}\|_{F}^{2}+2\|\mathbf{B}\|_{F}^{2}$:
-$\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\leq 2\beta_{t}^{2}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}^{2}$
+$\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\leq 2\beta_{t}^{2}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t})\|_{F}^{2}.$
 Thus, we have
@@
 $\leq 2\beta_{t}^{2}\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\sigma^{2}$
-$\leq 2\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\sigma^{2}$
+$\leq 2\mathbb{E}\|\mathbf{C}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+2(1-\beta_{t})^{2}\sigma^{2}.$
 Thus,
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}4\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)+2\sigma^{2}$
-$=6\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)$
+$=6\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1),$
 where $(\circ)$ is due to $\sum_{t=1}^{T}\frac{1}{t(t+1)}=\sum_{t=1}^{T}(\frac{1}{t}-\frac{1}{t+1})=1-\frac{1}{T+1}\leq 1$.
@@
 $\Gamma_{t}$ $=\frac{\eta_{t}^{2/3}}{L}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\eta_{t}^{4/3}$
-$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}$
+$=\frac{\alpha_{t}}{L}\mathbb{E}\|\mathbf{S}^{\prime}_{t}\|_{F}^{2}+(L/2+Ln)\alpha_{t}^{2}.$
 Then, we define $A_{1}=8\sqrt{2}Ln+Ln+4L^{-1}\sigma^{2}+L/2$ and $A_{2}=8\sqrt{2}Ln+Ln+10L^{-1}\sigma^{2}+L/2$. We have
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{6\sigma^{2}+4(2\sqrt{2}L^{2}n+\sigma^{2})(\ln T+1)}{L}+(L/2+Ln)(\ln T+1)$
-$\leq A_{1}\ln T+A_{2}$
+$\leq A_{1}\ln T+A_{2},$
 where ($\circ$) is due to inequality (12).
@@
 $\leq\frac{f(\mathbf{X}_{1})-f^{*}}{T^{1/4}}+\frac{A_{1}\ln T+A_{2}}{T^{1/4}}$
-$=\tilde{\mathcal{O}}(T^{-1/4})$
+$=\tilde{\mathcal{O}}(T^{-1/4}).$
 This completes the proof. ∎
@@
 $\stackrel{{\scriptstyle(\star)}}{{\geq}}\ln(t)t^{-q}\cdot\left\{1\cdot\sqrt{\frac{t}{t+1}}-\frac{1}{4}\right\}$
-$\stackrel{{\scriptstyle(\bullet)}}{{\geq}}\ln(t)t^{-q}\cdot\left\{\sqrt{\frac{1}{2}}-\frac{1}{4}\right\}\geq 0$
-Here, $(\circ)$ holds because $\ln(t+1)>\ln(t)$ and $(t+1)^{q}>t^{q}$ for all $t\geq 1$ and $q\in(0,1/2]$.
+$\stackrel{{\scriptstyle(\bullet)}}{{\geq}}\ln(t)t^{-q}\cdot\left\{\sqrt{\frac{1}{2}}-\frac{1}{4}\right\}\geq 0,$
+where $(\circ)$ holds because $\ln(t+1)>\ln(t)$ and $(t+1)^{q}>t^{q}$ for all $t\geq 1$ and $q\in(0,1/2]$.
 The inequality $(\star)$ holds because $\frac{\ln(t+1)}{\ln(t)}>1$ for $t>1$, and the function $(\frac{t}{t+1})^{q}$ is decreasing in $q$, thus its minimum on $(0,1/2]$ is achieved at $q=1/2$.
@@
 Finally, for all $t>1$ and $q\in(0,1/2]$, we have
-$F(t,q)\leq F(t,1/2)\leq 0$ (15)
+$F(t,q)\leq F(t,1/2)\leq 0,$ (15)
 where the first inequality holds because $f(q)$ is monotonically increasing in $q$ on $(0,1/2]$ for any fixed $t>1$.
@@
 $\stackrel{{\scriptstyle(\star)}}{{=}}4A_{t}\cdot\left(\frac{1}{4}t^{-q}-t^{q}+(t+1)^{q}-(t+1)^{-q}\right)$
-$\stackrel{{\scriptstyle(\bullet)}}{{\leq}}0$
+$\stackrel{{\scriptstyle(\bullet)}}{{\leq}}0.$
 Here, $(\circ)$ follows from the assumption $A_{t+1}\leq(1-\varepsilon_{t+1})A_{t}+B_{t+1}$.
@@
 Then we have the following inequality:
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}+P_{t+1}$
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}+P_{t+1},$
 where
@@
 $B_{t+1}$ $=(1-\beta_{t+1})\mathbb{E}\langle\Delta_{t},\mathbf{R}_{t+1}\rangle+\beta_{t+1}\mathbb{E}\langle\Delta_{t},\mathbf{S}_{t}\rangle$
-$P_{t+1}$ $=\mathbb{E}\|\Delta_{t}\|_{F}^{2}(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1})^{2}-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}$
+$P_{t+1}$ $=\mathbb{E}\|\Delta_{t}\|_{F}^{2}(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1})^{2}-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}.$
 If we choose $\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{t+1}}$ or $\gamma_{t+1}=1$, then
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}.$
 Proof.
@@
 $=(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\beta_{t+1}(\mathbf{M}_{t}+\gamma_{t+1}(\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t};\xi_{t+1})))$
-$=(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\beta_{t+1}(\mathbf{M}_{t}+\gamma_{t+1}\Delta_{t})$
+$=(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})+\beta_{t+1}(\mathbf{M}_{t}+\gamma_{t+1}\Delta_{t}).$
 Hence,
@@
 $\quad+\beta_{t+1}\left(\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})+\delta_{t}+\gamma_{t+1}\Delta_{t}\right)$
-$=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}(\mathbf{S}_{t}+\delta_{t}+\gamma_{t+1}\Delta_{t})$
+$=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}(\mathbf{S}_{t}+\delta_{t}+\gamma_{t+1}\Delta_{t}).$
 Note that, according to Assumption 3.4:
-$\delta_{t}=\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})=-\mathbb{E}_{t+1}[\Delta_{t}]$
+$\delta_{t}=\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})=-\mathbb{E}_{t+1}[\Delta_{t}].$
 Therefore,
-$\mathbf{S}_{t+1}=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\gamma_{t+1}\Delta_{t}-\mathbb{E}\Delta_{t})$
+$\mathbf{S}_{t+1}=(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\gamma_{t+1}\Delta_{t}-\mathbb{E}\Delta_{t}).$
 Thus,
@@
 $\quad+\underbrace{\beta_{t+1}^{2}(\gamma_{t+1}-1)^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}}_{\text{Term A.2}}$
-$\quad+\underbrace{2\beta_{t+1}(\gamma_{t+1}-1)\mathbb{E}\langle\Delta_{t},(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\rangle}_{\text{Term A.3}}$
+$\quad+\underbrace{2\beta_{t+1}(\gamma_{t+1}-1)\mathbb{E}\langle\Delta_{t},(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}\mathbf{S}_{t}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\rangle}_{\text{Term A.3}}.$
 First, let’s consider Term A.1:
@@
 $\stackrel{{\scriptstyle(\circ)}}{{=}}\mathbb{E}\|(1-\beta_{t+1})\mathbf{R}_{t+1}+\beta_{t+1}(\Delta_{t}-\mathbb{E}\Delta_{t})\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},$
 where $(\circ)$ holds because Assumption 3.4 implies $\mathbb{E}[\mathbf{R}_{t+1}]=0$ and $\mathbb{E}[\Delta_{t}-\mathbb{E}\Delta_{t}]=0$, making the cross-terms with $\mathbf{S}_{t}$ zero; and $(\star)$ follows from the inequality $\|\mathbf{A}+\mathbf{B}\|_{F}^{2}\leq 2\|\mathbf{A}\|_{F}^{2}+2\|\mathbf{B}\|_{F}^{2}$ for any $\mathbf{A},\mathbf{B}\in\mathbb{R}^{m\times n}$.
@@
 $A_{t+1}$ $=\frac{B_{t+1}+\beta_{t+1}(\mathbb{E}\|\Delta_{t}\|_{F}^{2}-\|\mathbb{E}\Delta_{t}\|_{F}^{2})}{\mathbb{E}\|\Delta_{t}\|_{F}^{2}}$
-$B_{t+1}$ $=(1-\beta_{t+1})\mathbb{E}\langle\Delta_{t},\mathbf{R}_{t+1}\rangle+\beta_{t+1}\mathbb{E}\langle\Delta_{t},\mathbf{S}_{t}\rangle$
+$B_{t+1}$ $=(1-\beta_{t+1})\mathbb{E}\langle\Delta_{t},\mathbf{R}_{t+1}\rangle+\beta_{t+1}\mathbb{E}\langle\Delta_{t},\mathbf{S}_{t}\rangle.$
 Then we have
@@
 $=\mathbb{E}\|\Delta_{t}\|_{F}^{2}(\beta_{t+1}(1-\gamma_{t+1})-A_{t+1})^{2}-A_{t+1}^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}$
-$=P_{t+1}$
+$=P_{t+1}.$
 Therefore, combining the bounds:
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}2(1-\beta_{t+1})^{2}\sigma^{2}+2\beta_{t+1}^{2}\mathbb{E}\|\Delta_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1}$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\sigma^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1}$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}2(1-\beta_{t+1})^{2}\sigma^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+P_{t+1},$
 where $(\circ)$ uses Assumption 3.4 (bounded variance, i.e., $\mathbb{E}\|\mathbf{R}_{t+1}\|_{F}^{2}\leq\sigma^{2}$) and the property that $\mathbb{E}\|\Delta_{t}-\mathbb{E}\Delta_{t}\|_{F}^{2}\leq\mathbb{E}\|\Delta_{t}\|_{F}^{2}$; and $(\star)$ follows from Assumption 3.3, which implies $\|\Delta_{t}\|_{F}^{2}\leq L^{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$.
 Next, we set
-$\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{t+1}}\text{ or }\gamma_{t+1}=\gamma=1$ (16)
+$\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{t+1}}\text{ or }\gamma_{t+1}=\gamma=1.$ (16)
 Then,
-$P_{t+1}=-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}\leq 0\text{ or }P_{t+1}=0$
+$P_{t+1}=-\mathbb{E}\|\Delta_{t}\|_{F}^{2}A_{t+1}^{2}\leq 0\text{ or }P_{t+1}=0.$
 This leads to the final result:
 $\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}$ $\leq\beta_{t+1}^{2}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$
-$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2}$
+$\stackrel{{\scriptstyle(\circ)}}{{\leq}}\beta_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\beta_{t+1}^{2}L^{2}\mathbb{E}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}+2(1-\beta_{t+1})^{2}\sigma^{2},$
 where $(\circ)$ follows from $\beta_{t+1}^{2}\leq\beta_{t+1}\leq 1$. ∎
@@
 The proof begins with the recursive inequality derived from a preceding Lemma C.2:
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t}^{2}L^{2}n+2\eta_{t+1}^{2}\sigma^{2}$
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t}^{2}L^{2}n+2\eta_{t+1}^{2}\sigma^{2},$
 where $\mathbf{S}_{t+1}=\nabla f(\mathbf{X}_{t+1})-\mathbf{M}_{t+1}$. We have noticed the following facts
-$\frac{1}{t^{2/3}}\leq\frac{2}{(t+1)^{2/3}}$
+$\frac{1}{t^{2/3}}\leq\frac{2}{(t+1)^{2/3}}.$
 Therefore , we have
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})$
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-\eta_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2}).$
 Let $A_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ and $B_{t+1}=2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})$.
@@
 This structure allows us to apply a standard result Lemma C.1 which yields:
-$\sqrt{\eta_{t}}A_{t}\leq 4\left(\frac{A_{t}}{\sqrt{\eta_{t}}}-\frac{A_{t+1}}{\sqrt{\eta_{t+1}}}+\frac{B_{t+1}}{\sqrt{\eta_{t+1}}}\right)$
+$\sqrt{\eta_{t}}A_{t}\leq 4\left(\frac{A_{t}}{\sqrt{\eta_{t}}}-\frac{A_{t+1}}{\sqrt{\eta_{t+1}}}+\frac{B_{t+1}}{\sqrt{\eta_{t+1}}}\right).$
 Then, we define $P_{t}=\frac{4A_{t}}{\sqrt{\eta_{t}}}=\frac{4\mathbb{E}\|\mathbf{S}_{t}\|^{2}_{F}}{\sqrt{\eta_{t}}}$. Substituting $P_{t}$ and the definition of $B_{t+1}$ into the inequality gives:
-$\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq P_{t}-P_{t+1}+\frac{4\cdot 2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})}{\sqrt{\eta_{t+1}}}=P_{t}-P_{t+1}+8\eta_{t+1}^{3/2}(2L^{2}n+\sigma^{2})$
+$\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq P_{t}-P_{t+1}+\frac{4\cdot 2\eta_{t+1}^{2}(2L^{2}n+\sigma^{2})}{\sqrt{\eta_{t+1}}}=P_{t}-P_{t+1}+8\eta_{t+1}^{3/2}(2L^{2}n+\sigma^{2}).$
 Now, we sum this inequality from $t=1$ to $T$:
 $\sum_{t=1}^{T}\sqrt{\eta_{t}}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq\sum_{t=1}^{T}\left(P_{t}-P_{t+1}+8\eta_{t+1}^{3/2}(2L^{2}n+\sigma^{2})\right)$
-$=(P_{1}-P_{T+1})+8(2L^{2}n+\sigma^{2})\sum_{t=1}^{T}\eta_{t+1}^{3/2}$
+$=(P_{1}-P_{T+1})+8(2L^{2}n+\sigma^{2})\sum_{t=1}^{T}\eta_{t+1}^{3/2}.$
 Since $P_{T+1}\geq 0$, we can drop this term to simplify the bound. By setting the step size $\eta_{t}=t^{-2/3}$, we have $\eta_{t+1}^{3/2}=((t+1)^{-2/3})^{3/2}=(t+1)^{-1}$. The summation becomes:
@@
 $\leq P_{1}+8(2L^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t}$
-$\stackrel{{\scriptstyle(\circ)}}{{\leq}}P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)$
+$\stackrel{{\scriptstyle(\circ)}}{{\leq}}P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T),$
 where $(\circ)$ follows from the harmonic series, $\sum_{t=1}^{T}\frac{1}{t}\leq 1+\ln T$.
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{T^{1/3}}{T}\left(P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)\right)$
-$=\frac{P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}$
+$=\frac{P_{1}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}},$
 where $(\circ)$ follows from the inequality (17).
@@
 $=4\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}4\sigma^{2}$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}4\sigma^{2}.$
 The final inequality $(\star)$ holds by setting the parameter $\gamma_{1}=\gamma=1$.
@@
 $\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{*}+\frac{1}{2\alpha}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\alpha+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2},$
 where $(\circ)$ holds because, by the definition of $\mathbf{O}_{t}$ and the property of norms $\|\mathbf{M}_{t}\|_{*}\geq\|\mathbf{M}_{t}\|_{F}$, we have:
-$\langle\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F}$
+$\langle\mathbf{M}_{t},\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=-\eta_{t}\langle\mathbf{M}_{t},\mathbf{U}_{r}\mathbf{V}_{r}^{\top}\rangle=-\eta_{t}\|\mathbf{M}_{t}\|_{*}\leq-\eta_{t}\|\mathbf{M}_{t}\|_{F},$
 and $(\star)$ holds by setting $\alpha=\frac{L}{\sqrt{\eta_{t}}}$. Thus, we have
@@
 $\quad+\frac{Ln}{2}(1+\ln T)+\frac{Ln}{2}\sum_{t=1}^{T}t^{-4/3}$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+2Ln+4(2Ln+\sigma^{2}L^{-1})(1+\ln T)+\frac{Ln}{2}(1+\ln T)$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+2Ln+4(2Ln+\sigma^{2}L^{-1})(1+\ln T)+\frac{Ln}{2}(1+\ln T),$
 where $(\circ)$ uses Lemma C.2; $(\star)$ follows from the fact that $\sum_{t=1}^{T}\frac{1}{t^{4/3}}\leq 4$. Next, we let
-$G=f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+\left(8Ln+4\sigma^{2}L^{-1}+\frac{Ln}{2}\right)(1+\ln T)+2Ln$
+$G=f(\mathbf{X}_{1})-f^{*}+\frac{2\sigma^{2}}{L}+\left(8Ln+4\sigma^{2}L^{-1}+\frac{Ln}{2}\right)(1+\ln T)+2Ln.$
 Thus, we have
@@
 $\leq\frac{T^{2/3}}{T}\sum_{t=1}^{T}\frac{1}{t^{2/3}}\mathbb{E}\|\mathbf{M}_{t}\|_{F}=\frac{1}{T^{1/3}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\mathbf{M}_{t}\|_{F}$
-$\leq\frac{G}{T^{1/3}}$
+$\leq\frac{G}{T^{1/3}}.$
 Next, we have
 $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}$ $\stackrel{{\scriptstyle(\circ)}}{{\leq}}\sqrt{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}}$
-$\stackrel{{\scriptstyle(\star)}}{{\leq}}\sqrt{\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}}$
+$\stackrel{{\scriptstyle(\star)}}{{\leq}}\sqrt{\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}}.$
 where $(\circ)$ uses Jensen’s inequality; $(\star)$ uses Lemma C.3 by letting $\mathbf{S}_{t}=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}$. Thus, we have
@@
 $\leq\sqrt{\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{T^{2/3}}}+\frac{G}{T^{1/3}}$
-$=\mathcal{O}\left(\frac{\ln T}{T^{1/3}}\right)$
+$=\mathcal{O}\left(\frac{\ln T}{T^{1/3}}\right).$
 This completes the proof. ∎
-Appendix E Lemmas for Theorems 4.3 and 4.4
+Appendix E Lemmas for Theorems 3.3 and 3.4
 E.1 Lemma E.1
@@
 Define the function $f(t)=\frac{t^{1-p}}{\ln t}$ where $p\in(0,1)$.
-Then, for any sufficiently large integer $T_{0}>e^{1/(1-p)}$ where $e$ is the base of the natural logarithm, the following inequality holds for all $T>T_{0}$:
-$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq\left(A_{1}+A_{2}\right)(T-1)^{1-p}$
+Fix $T_{0}=2e^{1/(1-p)}$, where $e$ is the base of the natural logarithm.
+Then, the following inequality holds for all $T>T_{0}$:
+$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq\left(A_{1}+A_{2}\right)(T-1)^{1-p}.$
 Proof.
-We choose an integer $T_{0}>e^{1/(1-p)}$ large enough such that for all $t\geq T_{0}$, the function $f(t)$ is positive and monotonically increasing.
+We choose an integer $T_{0}=2e^{1/(1-p)}>e^{1/(1-p)}$ large enough such that for all $t\geq T_{0}$, the function $f(t)$ is positive and monotonically increasing.
 We use summation by parts to bound the sum $\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)$:
-$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)=\sum_{t=T_{0}}^{T-1}(D_{t}-D_{t-1})f(t)=D_{T-1}f(T-1)-D_{T_{0}-1}f(T_{0})-\sum_{t=T_{0}}^{T-1}D_{t-1}(f(t)-f(t-1))$
+$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)=\sum_{t=T_{0}}^{T-1}(D_{t}-D_{t-1})f(t)=D_{T-1}f(T-1)-D_{T_{0}-1}f(T_{0})-\sum_{t=T_{0}}^{T-1}D_{t-1}(f(t)-f(t-1)).$
 Since $D_{t-1}\geq 0$ and $f(t)$ is monotonically increasing for $t\geq T_{0}$ (i.e., $f(t)-f(t-1)\geq 0$), the last term, $\sum_{t=T_{0}}^{T-1}D_{t-1}(f(t)-f(t-1))$, is non-negative.
 Therefore, we have:
-$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq D_{T-1}f(T-1)$
+$\sum_{t=T_{0}}^{T-1}\Gamma_{t}f(t)\leq D_{T-1}f(T-1).$
 Substituting the given bound on the partial sums, $D_{T-1}\leq A_{1}\ln(T-1)+A_{2}$, we obtain:
@@
 $=\left(A_{1}+\frac{A_{2}}{\ln(T-1)}\right)(T-1)^{1-p}$
-$\stackrel{{\scriptstyle(\circ)}}{{\leq}}(A_{1}+A_{2})(T-1)^{1-p}$
+$\stackrel{{\scriptstyle(\circ)}}{{\leq}}(A_{1}+A_{2})(T-1)^{1-p},$
 where $(\circ)$ follows from $\ln(T-1)\geq 1$ for sufficiently large $T$.
@@
 Suppose a positive sequence $\{\Delta_{t}\}_{t\geq 1}$ satisfies the following recursive inequality:
-$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}$
+$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t},$
 where $p\in(0,1)$ and $\mu>0$ are constants.
-The non-negative noise sequence $\{\Gamma_{t}\}_{t\geq 1}$ satisfies the condition from Lemma E.1, with its partial sum bounded by $\sum_{t=1}^{T}\Gamma_{t}=\mathcal{O}(\ln T)$.
+The non-negative noise sequence $\{\Gamma_{t}\}_{t\geq 1}$ satisfies the condition from Lemma E.1 that $\sum_{t=1}^{T}\Gamma_{t}=A_{1}\ln T+A_{2}$.
 Then, the sequence $\Delta_{t}$ converges to 0 with the following rate:
-$\Delta_{t}=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right)$
+$\Delta_{t}\leq\frac{\mathcal{A}^{2}(1-p)^{2}}{C^{2}}\cdot\frac{(\ln t)^{2}}{t^{2(1-p)}},$
+where $\mathcal{A}:=A_{1}+A_{2}$ and $C=\frac{\sqrt{2\mu}}{2}$.
 Proof.
-We first define the variable $G_{t}=\sqrt{\Delta_{t}}$.
+We define the variable $G_{t}=\sqrt{\Delta_{t}}$.
 Since $\Delta_{t}>0$, it follows that $G_{t}>0$.
@@
 By setting $y=\Delta_{t+1}$ and $x=\Delta_{t}$, we obtain:
-$G_{t+1}-G_{t}\leq\frac{\Delta_{t+1}-\Delta_{t}}{2G_{t}}$
+$G_{t+1}-G_{t}\leq\frac{\Delta_{t+1}-\Delta_{t}}{2G_{t}}.$
 Substituting the original recurrence relation $\Delta_{t+1}-\Delta_{t}\leq-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}$ and noting that $\sqrt{\Delta_{t}}=G_{t}$, we have:
-$G_{t+1}-G_{t}\leq\frac{1}{2G_{t}}\left(-\frac{\sqrt{2\mu}}{t^{p}}G_{t}+\Gamma_{t}\right)$
-Letting $C=\frac{\sqrt{2\mu}}{2}$, we arrive at the core recurrence for $G_{t}$:
+$G_{t+1}-G_{t}\leq\frac{1}{2G_{t}}\left(-\frac{\sqrt{2\mu}}{t^{p}}G_{t}+\Gamma_{t}\right).$
+Letting $C=\frac{\sqrt{2\mu}}{2},T_{0}=2e^{1/(1-p)}$, we arrive at the core recurrence for $G_{t}$:
 $G_{t+1}\leq G_{t}-\frac{C}{t^{p}}+\frac{\Gamma_{t}}{2G_{t}}$ (18)
-For $p\in(0,1)$, we will prove that $G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1-p}}\right)$ by contradiction.
+For $p\in(0,1)$, we will prove that $G_{t}\leq\frac{\mathcal{A}(1-p)}{C}\cdot\frac{\ln t}{t^{1-p}}$ by contradiction.
 Assume the proposition is false. This means that for any constant $A>0$, there exist infinitely many time steps $t$ such that $G_{t}>B_{t}$, where $B_{t}=A\cdot\frac{\ln t}{t^{1-p}}$.
@@
 Summing the inequality (18) from $t=T_{0}$ to T1T-1 gives:
-$G_{T}-G_{T_{0}}\leq\sum_{t=T_{0}}^{T-1}(G_{t+1}-G_{t})\leq\sum_{t=T_{0}}^{T-1}\left(-\frac{C}{t^{p}}+\frac{\Gamma_{t}}{2G_{t}}\right)$
+$G_{T}-G_{T_{0}}\leq\sum_{t=T_{0}}^{T-1}(G_{t+1}-G_{t})\leq\sum_{t=T_{0}}^{T-1}\left(-\frac{C}{t^{p}}+\frac{\Gamma_{t}}{2G_{t}}\right).$
 Over this interval, since $G_{t}>B_{t}=A\frac{\ln t}{t^{1-p}}$, we have $\frac{1}{G_{t}}<\frac{1}{B_{t}}=\frac{t^{1-p}}{A\ln t}$. Substituting this bound yields:
-$G_{T}<G_{T_{0}}-\underbrace{C\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}}_{V_{1}}+\underbrace{\frac{1}{2A}\sum_{t=T_{0}}^{T-1}\frac{\Gamma_{t}t^{1-p}}{\ln t}}_{V_{2}}$
+$G_{T}<G_{T_{0}}-\underbrace{C\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}}_{V_{1}}+\underbrace{\frac{1}{2A}\sum_{t=T_{0}}^{T-1}\frac{\Gamma_{t}t^{1-p}}{\ln t}}_{V_{2}}.$
 Next, we analyze the asymptotic behavior of the two sums, $V_{1}$ and $V_{2}$.
 For the negative drift term $V_{1}$, we use an integral approximation:
-$\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}\geq\int_{T_{0}}^{T}x^{-p}dx=\frac{1}{1-p}(T^{1-p}-T_{0}^{1-p})$
+$\sum_{t=T_{0}}^{T-1}\frac{1}{t^{p}}\geq\int_{T_{0}}^{T}x^{-p}dx=\frac{1}{1-p}(T^{1-p}-T_{0}^{1-p}).$
 The leading term of $V_{1}$ is thus on the order of $\frac{C}{1-p}T^{1-p}$.
@@
 Substituting the leading terms of these bounds back into the inequality for $G_{T}$:
-$G_{T}<G_{T_{0}}-\frac{C}{1-p}T^{1-p}+\frac{\mathcal{A}}{2A}T^{1-p}=G_{T_{0}}+T^{1-p}\left(\frac{\mathcal{A}}{2A}-\frac{C}{1-p}\right)$
+$G_{T}<G_{T_{0}}-\frac{C}{1-p}T^{1-p}+\frac{\mathcal{A}}{2A}T^{1-p}=G_{T_{0}}+T^{1-p}\left(\frac{\mathcal{A}}{2A}-\frac{C}{1-p}\right).$
 We now choose the constant $A$ such that the coefficient of the leading term is negative:
-$\frac{\mathcal{A}}{2A}-\frac{C}{1-p}<0\implies A>\frac{\mathcal{A}}{2C}$
+$\frac{\mathcal{A}}{2A}-\frac{C}{1-p}<0\implies A>\frac{\mathcal{A}(1-p)}{2C}.$
 With an $A$ satisfying this condition, for a sufficiently large $T$, the negative term proportional to $T^{1-p}$ will dominate, forcing the right-hand side to become negative.
@@
 This contradiction proves that our assumption that $G_{t}$ can remain above $B_{t}$ for an arbitrarily long interval is false.
-Therefore, there must exist a time $T_{A}$ such that for all $t>T_{A}>T_{0}$, $G_{t}\leq B_{t}$.
+Therefore, there must exist a time $T_{A}$ such that for all $t>T_{A}>T_{0}=2e^{1/(1-p)}$, $G_{t}\leq B_{t}$.
 This establishes that $G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1-p}}\right)$.
-Finally, squaring both sides of this result yields the convergence rate for $\Delta_{t}$:
-$\Delta_{t}=G_{t}^{2}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1-p}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right)$
+Finally, by choosing the constant $A=\frac{\mathcal{A}(1-p)}{C}$ , squaring both sides of this result yields the convergence rate for $\Delta_{t}$:
+$\Delta_{t}=G_{t}^{2}$ $\leq\frac{\mathcal{A}^{2}(1-p)^{2}}{C^{2}}\cdot\frac{(\ln t)^{2}}{t^{2(1-p)}}$
+$=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2(1-p)}}\right).$
 Appendix F Proofs of Theorem 3.3
@@
 Let $\eta_{t}=t^{-3/4},\beta_{t}=1-t^{-1/2},\alpha_{t}=t^{-1/2}$. According to Theorem 3.1 inequality (9), we have
-$\sum_{t=1}^{T}\Gamma_{t}\leq A_{1}\ln T+A_{2}$
+$\sum_{t=1}^{T}\Gamma_{t}\leq A_{1}\ln T+A_{2},$
 where $A_{1}=2L^{-1}\sigma^{2}+4\sqrt{2}Ln+Ln+L/2$ and $A_{2}=4L^{-1}\sigma^{2}+4\sqrt{2}Ln+Ln+L/2$.
@@
 Let $\eta_{t}=t^{-3/4},\beta_{t}=1-(1+t)^{-1/2},\alpha_{t}=t^{-1/2}$. According to Theorem 3.1 inequality (13), we have
-$\sum_{t=1}^{T}\Gamma_{t}\leq A_{1}\ln T+A_{2}$
+$\sum_{t=1}^{T}\Gamma_{t}\leq A_{1}\ln T+A_{2},$
 where $A_{1}=4L^{-1}\sigma^{2}+8\sqrt{2}Ln+Ln+L/2$ and $A_{2}=10L^{-1}\sigma^{2}+8\sqrt{2}Ln+Ln+L/2$.
@@
 $\Delta_{t+1}$ $\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\Delta_{t}}+\Gamma_{t}$
-$=\Delta_{t}-\frac{\sqrt{2\mu}}{t^{3/4}}\sqrt{\Delta_{t}}+\Gamma_{t}$
-By Lemma E.2, we have
-$G_{t+1}\leq G_{t}-\frac{C}{t^{3/4}}+\frac{\Gamma_{t}}{2G_{t}}$
+$=\Delta_{t}-\frac{\sqrt{2\mu}}{t^{3/4}}\sqrt{\Delta_{t}}+\Gamma_{t}.$
+By Lemma E.2, let $G_{t}=\sqrt{\Delta_{t}}$, $C=\sqrt{2\mu}/2$ and $p=3/4$.
+$G_{t+1}\leq G_{t}-\frac{C}{t^{3/4}}+\frac{\Gamma_{t}}{2G_{t}}.$
 Setting $p=3/4$, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{8C}$, there exists a time step $T_{A}$ such that for all $t>T_{A}$, we have $G_{t}\leq A\cdot\frac{\ln t}{t^{1/4}}$.
-Thus, we obtain the convergence rate for $G_{t}$:
-$G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1/4}}\right)$
-Finally, since $\Delta_{t}=G_{t}^{2}$, squaring both sides yields the final convergence rate of $\Delta_{t}$:
-$\Delta_{t}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1/4}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{1/2}}\right)$
-Hence,
-$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right)$
+By choosing $A=\frac{\mathcal{A}}{4C}$, we satisfy the condition $A>\frac{\mathcal{A}}{8C}$, which ensures this bound holds.
+Thus, we obtain the convergence rate for $\Delta_{t}$:
+$\Delta_{t}\leq\frac{\mathcal{A}^{2}(1-3/4)^{2}}{C^{2}}\cdot\frac{(\ln t)^{2}}{t^{2(1-3/4)}}=\frac{\mathcal{A}^{2}}{8\mu}\cdot\frac{(\ln t)^{2}}{t^{1/2}}$
+We can now summarize the results for both scenarios.
+In either case ($\gamma=0$ or $\gamma\neq 0$), the analysis yields the same asymptotic convergence rate.
+For any number of iterations $T>T_{0}=2e^{4}$, the analysis leads to the same asymptotic upper bound:
+$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{\mathcal{A}^{2}}{8\mu}\cdot\frac{(\ln T)^{2}}{T^{1/2}}=\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right).$
+The specific definition of $\mathcal{A}$ is as follows: (i) Case 1 ($\gamma=0$): $\mathcal{A}=6L^{-1}\sigma^{2}+8\sqrt{2}Ln+2Ln+L$.
+(ii) Case 2 ($\gamma\neq 0$): $\mathcal{A}=14L^{-1}\sigma^{2}+16\sqrt{2}Ln+2Ln+L$.
 This completes the proof. ∎
@@
 According to Theorem 3.2, we have
-$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
+$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\eta_{t}\|\mathbf{M}_{t}\|_{F}+\frac{\sqrt{\eta_{t}}}{2L}\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}+\frac{\frac{L}{\sqrt{\eta_{t}}}+L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}.$
 Since $\eta_{t}\leq 1$, we have $\frac{L}{\sqrt{\eta_{t}}}\geq L$. Thus, taking the expectation yields
@@
 $\stackrel{{\scriptstyle(\bullet)}}{{=}}\mathbb{E}[f(\mathbf{X}_{t})]-\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})\|_{F}]$
-$+\underbrace{\frac{L\eta_{t}^{3/2}}{2}+\left(\frac{\epsilon}{2}+\frac{\sqrt{\eta_{t}}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}}_{\Gamma_{t}}$
+$+\underbrace{\frac{L\eta_{t}^{3/2}}{2}+\left(\frac{\epsilon}{2}+\frac{\sqrt{\eta_{t}}}{2L}\right)\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}^{2}]+Ln\eta_{t}^{3/2}}_{\Gamma_{t}},$
 where $(\circ)$ follows from the reverse triangle inequality $\|\mathbf{M}_{t}\|_{F}\leq\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}-\|\nabla f(\mathbf{X}_{t})\|_{F}$; $(\star)$ applies Young’s inequality to the term $\eta_{t}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}\|_{F}]$; and $(\bullet)$ collects the residual terms into $\Gamma_{t}$.
 By setting $\epsilon=\frac{\sqrt{\eta_{t}}}{L}$, we can simplify the expression for $\Gamma_{t}$:
-$\Gamma_{t}=\frac{\sqrt{\eta_{t}}}{L}\mathbb{E}[\|\mathbf{S}_{t}\|_{F}^{2}]+(L/2+Ln)\eta_{t}^{3/2}$
+$\Gamma_{t}=\frac{\sqrt{\eta_{t}}}{L}\mathbb{E}[\|\mathbf{S}_{t}\|_{F}^{2}]+(L/2+Ln)\eta_{t}^{3/2},$
 where $\mathbf{S}_{t}=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t}$. The sum of these terms can be bounded. Let $A_{1}=17Ln+8L^{-1}\sigma^{2}+L/2$ and $A_{2}=17Ln+12L^{-1}\sigma^{2}+L/2$. We have
@@
 $\stackrel{{\scriptstyle(\circ)}}{{\leq}}\frac{4\sigma^{2}+8(2L^{2}n+\sigma^{2})(1+\ln T)}{L}+(L/2+Ln)(1+\ln T)$
-$=A_{1}\ln T+A_{2}$
+$=A_{1}\ln T+A_{2},$
 where $(\circ)$ holds by Lemma C.3 inequality (17).
@@
 $\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\mathbb{E}[f(\mathbf{X}_{t})-f^{*}]}+\Gamma_{t}$
-$\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\Delta_{t}}+\Gamma_{t}$
+$\leq\Delta_{t}-\eta_{t}\sqrt{2\mu\Delta_{t}}+\Gamma_{t}.$
 Setting the step size $\eta_{t}=1/t^{2/3}$, the recurrence becomes
-$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{2/3}}\sqrt{\Delta_{t}}+\Gamma_{t}$
+$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{2/3}}\sqrt{\Delta_{t}}+\Gamma_{t}.$
 This recurrence is in the form required by Lemma E.2.
@@
 According to Lemma E.2, for any constant $A>\frac{(A_{1}+A_{2})(1-p)}{2C}=\frac{A_{1}+A_{2}}{6C}$, there exists a time $T_{A}$ such that for all $t>T_{A}$, $G_{t}\leq A\cdot\frac{\ln t}{t^{1/3}}$.
-This establishes the convergence rate for $G_{t}$:
-$G_{t}=\mathcal{O}\left(\frac{\ln t}{t^{1/3}}\right)$
-Finally, since $\Delta_{t}=G_{t}^{2}$, squaring both sides gives the final convergence rate for $\Delta_{t}$:
-$\Delta_{t}=\mathcal{O}\left(\left(\frac{\ln t}{t^{1/3}}\right)^{2}\right)=\mathcal{O}\left(\frac{(\ln t)^{2}}{t^{2/3}}\right)$
-Therefore, we have
-$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right)$
+We choose a specific constant $A=\frac{\mathcal{A}}{3C}$, where $\mathcal{A}:=A_{1}+A_{2}=20L^{-1}\sigma^{2}+34Ln+L$.
+Squaring both sides of the inequality for $G_{t}$ establishes the convergence rate for $\Delta_{t}=G_{t}^{2}$:
+$\Delta_{t}\leq A^{2}\cdot\frac{(\ln t)^{2}}{t^{2/3}}=\frac{\mathcal{A}^{2}}{9C^{2}}\frac{(\ln t)^{2}}{t^{2/3}}=\frac{2\mathcal{A}^{2}}{9\mu}\frac{(\ln t)^{2}}{t^{2/3}}$
+Therefore, for any number of iterations $T>T_{0}=2e^{3}$, the expected suboptimality has the following asymptotic bound:
+$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{2\mathcal{A}^{2}}{9\mu}\cdot\frac{(\ln T)^{2}}{T^{2/3}}=\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right).$
 This completes the proof. ∎
@@
 H.1 Training on CIFAR10
-The ResNet18 model He et al.
-(2016) undergoes pretraining on the CIFAR-10 dataset with comprehensive hyperparameter specifications provided in Table 2.
+The ResNet18 model he2016deep undergoes pretraining on the CIFAR-10 dataset with comprehensive hyperparameter specifications provided in Table 2.
 For each optimizer, the learning rate is selected via a grid search over the set $\{1\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}\}$.
@@
 Table 2: Hyperparameters used for training ResNet18 on CIFAR10
-SGD Adam Muon-Type
+SGD Adam Muon-Variant
 Model Size 42.7M
@@
 H.2 Pretraining on C4
-The LLaMA2-130M modelTouvron et al.
-(2023) was pre-trained on the C4 (Colossal Clean Crawled Corpus) dataset Raffel et al.
-(2020).
+The LLaMA2-130M modelTouvron2023Llama2O was pre-trained on the C4 (Colossal Clean Crawled Corpus) dataset raffel2020exploring.
 Identical learning rates and scheduling protocols were systematically implemented across all optimizers during the training process.
@@
 Table 3: Hyperparameters used for training LLaMA2-130M on C4
-AdamW Muon-Type
+AdamW Muon-Variant
 Model Size 130M